### v1

# VULCA-Bench: A Multicultural Vision-Language Benchmark for Evaluating Cultural Understanding

Abstract
We introduce Vulca-Bench, a multicultural art-critique benchmark for evaluating Vision–Language Models’ (VLMs) cultural understanding beyond surface-level visual perception. Existing VLM benchmarks predominantly measure L1–L2 capabilities (object recognition, scene description, and factual question answering) while under-evaluate higher-order cultural interpretation. Vulca-Bench contains 7,410 matched image–critique pairs spanning eight cultural traditions, with Chinese–English bilingual coverage. We operationalise cultural understanding using a five-layer framework (L1–L5, from Visual Perception to Philosophical Aesthetics), instantiated as 225 culture-specific dimensions and supported by expert-written bilingual critiques. Our pilot results indicate that higher-layer reasoning (L3–L5) is consistently more challenging than visual and technical analysis (L1–L2). The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 in the supplementary materials.
VULCA-Bench: A Multicultural Vision-Language Benchmark
for Evaluating Cultural Understanding
Haorui Yu and Ramon Ruiz-Dolz and Diji Yang and Hang He and Fengrui Zhang and Qiufeng Yi
1 Introduction
Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks, including 90%+ on object detection Li et al. (2023) and 77.2% on VQAv2 Goyal et al. (2017). However, these benchmarks predominantly assess L1–L2 visual perception, which involves recognising objects, describing scenes, and answering factual questions. They fail to evaluate cultural understanding, namely the capacity to interpret symbolic meanings, appreciate aesthetic traditions, and engage with philosophical concepts embedded in visual content. Recent studies confirm this gap: GPT-4o achieves only 54.1% accuracy on Chinese traditional painting Zeng et al. (2024), and cross-cultural evaluations reveal significant Western bias in VLM performance Liu et al. (2025).
We define cultural understanding using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics). Higher layers require progressively deeper cultural knowledge (Section 3 provides full definitions). Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3) Zhang (2024), the artist’s lineage within the “Four Gentlemen” tradition (L4), or the aesthetic principle of 气韵生动 (qiyun shengdong, “spirit resonance”) that defines Chinese painting philosophy (L5) Jing (2023). This example illustrates that cultural understanding is not a single capability but a progression across interpretive levels, a progression that existing benchmarks fail to measure.
Recent cultural AI datasets (CulturalBench Chiu et al. (2025), CultureAtlas Fung et al. (2024), GIMMICK Schneider et al. (2025)) begin to address this gap. Initial work on Chinese painting critique Yu et al. (2025a) introduced quantitative VLM evaluation using 163 expert commentaries, revealing significant VLM-expert divergence. Complementary probing Yu et al. (2025b) found VLMs rely on “symbolic shortcuts,” excelling on Western festivals but failing on underrepresented traditions. However, these studies focused on single cultures; systematic cross-cultural evaluation remains absent. Existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
We introduce Vulca-Bench (Vision-Understanding-Language-Culture Assessment), a comprehensive multicultural art critique benchmark addressing these gaps. Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and 98% cultural-fact accuracy, operationalised through our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size). All critiques meet expert-quality standards (Chinese $\geq$150 characters, English $\geq$100 words) with cultural specialist validation. The complete dataset is available under CC BY 4.0 in the supplementary materials.
Vulca-Bench enables systematic evaluation of VLMs’ hierarchical cultural understanding, identification of L1–L5 performance gaps, equal-weighted cultural fairness probing on the balanced-pilot subset (N=336, 7 cultures), and development of culturally grounded VLM architectures. Figure 1 showcases representative artworks from each tradition.
This paper makes three contributions. (1) We present Vulca-Bench, a multicultural art-critique benchmark enabling evaluation of cultural interpretation beyond surface-level perception, comprising 7,410 image–critique pairs across eight traditions with 225 expert-defined dimensions. (2) We formalise the Cultural Symmetry Principle, which enforces schema and protocol parity across cultures (without requiring equal sample sizes), thereby supporting fair cross-cultural evaluation. (3) We provide pilot experiments and error analyses showing that Vulca-Bench exposes systematic failures in higher-layer cultural reasoning (L3–L5) that are not captured by standard VLM benchmarks.
Figure 1: Cross-cultural Case Gallery from Vulca-Bench (All 8 Traditions). Each artwork represents the highest L3–L5 dimension coverage for its cultural tradition: Chinese (16), Western (17), Japanese (14), Korean (13), Islamic (16), Indian (16), Mural (14), Hermitage (16). The full corpus contains 7,410 image–critique pairs across 8 traditions with 225 culture-specific dimensions.
2 Related Work
Standard VLM benchmarks focus on L1 perception: MME Fu et al. (2023) and SEED-Bench-2 Li et al. (2024) achieve 90%+ on object detection but lack cultural evaluation; POPE Li et al. (2023) tests object existence without symbolic significance; and specialised benchmarks such as MathVista Lu et al. (2024), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) evaluate domain reasoning but not cultural understanding. Cultural AI datasets have begun to address this gap, yet remain limited. CulturalBench Chiu et al. (2025) and CulturalVQA Nayak et al. (2024) adopt QA formats rather than generative critique; GIMMICK Schneider et al. (2025) reveals Western biases across 144 countries but employs recognition tasks; WuMKG Wan et al. (2024) provides a Chinese painting knowledge graph (104K entities) for retrieval, not generative evaluation; and CulTi Yuan et al. (2025) offers image-text pairs for heritage retrieval without hierarchical cultural dimensions.
Art and aesthetics datasets similarly fall short. WikiArt Saleh and Elgammal (2015) (77K paintings, 27 styles) and OmniArt Strezoski and Worring (2018) (429K artworks) enable style classification but lack expert critiques or cultural dimension annotations; ArtEmis Achlioptas et al. (2021) focuses on affective responses; and Zhang et al. Zhang (2024) model Chinese painting critique via Xie He’s Six Principles at L1–L2 only. Prior VLM art critique work Yu et al. (2025a, b) focused on single cultures or specific visual domains, leaving critical gaps: no hierarchical L1–L5 schema operationalised across multiple cultures, no non-Western aesthetic frameworks such as Chinese yijing or Japanese wabi-sabi, and no expert-quality generative critique data with cultural dimension coverage.
Vulca-Bench addresses all three gaps through cross-cultural extension to eight traditions. Unlike QA-based cultural benchmarks that test factual recall, or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5). Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (layer-wise drop analysis, dimension coverage scoring) unavailable in existing datasets.
3 Dataset Construction
3.1 Cultural Symmetry Principle
Vulca-Bench adopts the Cultural Symmetry Principle, defined as schema and protocol parity across cultures, rather than sample-size parity. Concretely, each cultural tradition is annotated using: (i) the same five-layer structure (L1–L5), (ii) a uniform annotation protocol, and (iii) the same quality thresholds. This design reduces Western-centric methodological bias while acknowledging that museum availability and expert access naturally produce imbalanced sample sizes.
Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage). We recommend using balanced variants for per-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
We operationalise Cultural Symmetry through three pillars. Framework symmetry ensures each culture adapts L1–L5 to its own aesthetic theory (e.g., rasa for Indian, wabi-sabi for Japanese), with culture-specific dimension counts (25–30) reflecting tradition complexity rather than imposing uniform categories. Annotation symmetry requires native expert annotators for each tradition, with bilingual critiques preserving cultural terminology through romanisation. Quality symmetry applies uniform thresholds equally across all cultures ($\geq$70% dimension coverage, $\geq$150 Chinese characters or 100 English words).
The 70% threshold derives from pilot annotator calibration: critiques covering fewer than 70% of dimensions typically lacked substantive L3–L5 engagement. We therefore set this threshold empirically, based on the observation that lower coverage correlated strongly with missing philosophical or historical analysis. Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates, distributed proportionally across cultures).
Beyond structural parity, we validate measurement alignment using three criteria. First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures; the resulting five-point layer profiles are highly similar (mean pairwise Pearson correlation across cultures $r=0.96$). Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.79$ for L1–L2 and $0.71$ for L3–L5, with cross-culture standard deviation below 0.04. Third, each culture’s L5 captures its deepest interpretive challenge, supporting functional equivalence. See Appendix B for a finer-grained description of the corpus (i.e., complete, evaluation, gold, human, and balanced splits).
3.2 Cultural Understanding Framework
We operationalise cultural understanding through our five-layer framework reflecting increasing interpretive complexity. L1–L2 require visual observation; L3–L5 demand progressively deeper cultural knowledge that VLMs typically lack.
Layer Description
L1 Visual Perception: Color palette, line quality, composition, spatial layout, brushwork
L2 Technical Analysis: Medium, materials, craftsmanship, preservation state
L3 Cultural Symbolism: Motifs, iconography, narrative, symbolic meanings
L4 Historical Context: Period, artist biography, schools, provenance, influences
L5 Philosophical Aesthetics: Artistic conception, aesthetic theory, cultural values, innovation
Table 1: Five-layer cultural understanding framework (L1–L5). Higher layers require deeper cultural knowledge and interpretive reasoning.
Each culture adapts L1–L5 to its indigenous aesthetic vocabulary, yielding 225 dimensions total. Dimension counts (25–30) reflect tradition complexity: Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境); Western 25 dimensions emphasise formal analysis (chiaroscuro, perspective). All critiques must cover $\geq$70% of culture-specific dimensions, ensuring substantive L3–L5 engagement.
Culture Dims Pairs Key Concepts
Western 25 4,041 Chiaroscuro, linear perspective, impasto
Chinese 30 2,042 气韵 (qiyun), 笔墨 (bimo), 意境 (yijing)
Japanese 27 401 Wabi-sabi, yugen, ma (negative space)
Islamic 28 205 Geometric patterns, arabesques, Persian miniature
Mural 30 201 Buddhist art, Dunhuang murals, cave painting
Hermitage 30 196 Russian/European court art, iconography
Indian 30 173 Rasa theory (9 emotions), shringara, bhakti
Korean 25 151 Literati aesthetics, muninhwa tradition
Total 225 7,410 8 cultures, Cultural Symmetry
Table 2: Culture-specific dimensions in Vulca-Bench. Dimension counts reflect each tradition’s aesthetic complexity; threshold is $\geq$70% coverage per critique.
3.3 Data Sources
We curate artworks from authoritative museum collections with open-access policies. Chinese artworks derive from the Palace Museum, Shanghai Museum, and Metropolitan Museum; Western works from the Metropolitan Museum, Louvre, and Hermitage Museum; Japanese from Tokyo and Kyoto National Museums; Korean from the National Museum of Korea; Islamic from the Museum of Islamic Art Doha and Metropolitan Museum; Indian from the National Museum Delhi and British Museum; and Hermitage works from the State Hermitage Museum collection (Russian and European court art, religious iconography). All images are released under CC BY 4.0 in the supplementary materials.
3.4 Bilingual Critique Annotation
We adopt Chinese–English bilingual critiques for three reasons. First, terminology preservation: many Chinese aesthetic concepts (e.g., qiyun 气韵 “spirit resonance,” yijing 意境 “artistic conception”) are more precisely expressed in Chinese and are critical for consistent dimension annotation and automated matching. Second, international accessibility: English enables broader scholarly use while retaining romanised culture-specific terms where appropriate. Third, expert availability: our annotation team (Table 3) comprises bilingual specialists with deep cultural expertise, enabling controlled drafting, translation, and review within a single protocol. This design reflects practical constraints rather than linguistic preference; extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in Limitations.
Culture Ann. Rev. Background
Chinese 3 1 PhD art hist.; native CN
Western 2 1 PhD art hist.; EN native
Japanese 2 1 MA+ Japanese art; JP/CN
Korean 1 1 PhD Korean studies
Islamic 2 1 PhD Islamic art; AR/EN
Indian 2 1 PhD South Asian art
Table 3: Annotation team by culture. Ann.=annotators, Rev.=reviewers. All have 10+ years specialisation. Reviewers resolve disagreements and ensure cultural accuracy.
Expert annotators generate critiques following the L1–L5 framework. Table 3 summarises the annotation team. Each annotator receives the artwork image and metadata (title, artist, period), then generates a critique covering all 5 layers with culture-specific dimensions. Chinese critiques must reach $\geq$150 characters (average 450), and English translations preserve cultural terminology with romanisation. A reviewer validates each critique on dimension coverage ($\geq$70%), cultural accuracy, and bilingual consistency. Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources; unresolved cases (<<2%) escalate to senior expert panel. Full annotation guidelines appear in Appendix E; representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
We measure inter-annotator agreement on a 100-sample subset (Table 4). Cohen’s $\kappa=0.77$ (substantial agreement) for binary dimension presence; intraclass correlation coefficient ICC(2,1)=0.85=0.85 for dimension count. L3–L5 shows lower but still substantial agreement ($\kappa=0.72$), reflecting inherent interpretive complexity. Translation consistency achieved 94% agreement on cultural terminology preservation.
Korean follows a modified protocol due to the scarcity of qualified Korean art-history specialists with dual fluency. Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with 100% expert review by a senior scholar. The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA; we report Korean separately for methodological transparency.
Culture N L1–L2 $\kappa$ L3–L5 $\kappa$ ICC
Chinese 30 0.82 0.73 0.86
Western 25 0.79 0.69 0.83
Japanese 15 0.81 0.74 0.85
Islamic 10 0.77 0.72 0.84
Korean 10 0.75 0.70 0.82
Indian 10 0.76 0.71 0.83
Pooled 90 0.80 0.72 0.85
Table 4: Inter-annotator agreement by culture and layer. L3–L5 shows lower but substantial agreement, reflecting interpretive complexity. Pooled excludes Korean (5 cultures, N=90). Korean uses a different protocol: 1 primary annotator (PhD Korean studies) + 100% expert review. The reported $\kappa$ is annotator–reviewer agreement, not independent dual-annotation IAA; we report it separately for quality transparency but do not include it in pooled statistics.
3.5 Quality Assurance
Multi-stage validation ensures data quality. Each record contains an explicit dimension checklist through its covered_dimensions field, listing dimension IDs (e.g., ["CN_L1_D1", "CN_L3_D2", ...]) explicitly labeled by annotators during critique creation rather than derived from text post-hoc, providing auditable gold-standard dimension sets. Automated checks enforce the $\geq$70% dimension coverage threshold (21/30 for CN, 18/25 for WE, etc.) based on covered_dimensions length. We removed all exact duplicates after deep audit. Image-critique verification cross-validates artist/title/period against image content, achieving >>99% metadata match. All pairs have both Chinese ($\geq$150 characters) and English critiques, ensuring bilingual completeness. Finally, all images are compressed to $\leq$3.75MB for VLM API compatibility.
4 Dataset Statistics
Vulca-Bench comprises 7,410 matched image-critique pairs drawn from 11,964 total images across 8 cultural traditions, annotated with 225 culture-specific dimensions. Table 5 summarises key metrics: average Chinese critique length is 450 characters (English: 280 words), with 100% bilingual completeness, zero duplicates, >99% metadata verification, and 98% cultural-fact accuracy. All pairs meet the $\geq$70% dimension coverage threshold. Western (54.5%) and Chinese (27.6%) dominate, while minority cultures maintain rigorous verification. Audit levels range from full manual review (Korean) to stratified sampling (Indian).
Culture Pairs % Audit
Western 4,041 54.5% Deep
Chinese 2,042 27.6% Full
Japanese 401 5.4% Full
Islamic 205 2.8% Manual
Mural 201 2.7% Full
Hermitage 196 2.6% Full
Indian 173 2.3% Sample
Korean 151 2.0% Full manual
Total 7,410 100% Mixed
Zh: 450 chars avg; En: 280 words avg; 100% bilingual; 225 dims
[…]
Figure 2 reveals layer-wise coverage patterns: L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), while L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding. Chinese art shows highest L5 coverage (78%), consistent with its rich philosophical aesthetics tradition (qiyun, yijing). Importantly, each culture’s critiques are authored by native expert annotators from that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapanese
Figure 2: Layer-wise dimension coverage. L1–L2 achieve $\geq$94%; L3–L5 decline to 72–89%, reflecting interpretive complexity.
5 Data Quality Validation
We validate Vulca-Bench through a three-phase audit protocol. Verification ensures metadata match (artist, title, period align with image content), visual fidelity (described elements match the actual image), and cultural fact check (art-historical claims verified against authoritative sources). Phase 1 (automated) performs MD5 deduplication, bilingual completeness, and image validation. Phase 2 (stratified) constitutes the 450 human-scored subset. KR: 107 (full set), JP/IS/IN: 50 each, CN/WE: 100 each. Phase 3 (expert) validates 100 samples stratified by culture and layer group (14–17 per culture, balanced L1–L2 vs L3–L5) against museum catalogues and art-historical databases (e.g., Grove Art Online, Oxford Art Online). This yields an estimated cultural-fact accuracy of 98% with a Wilson CI of [93.0%, 99.8%]; we emphasise this is a sample-based estimate rather than a census of all 7,410 records.
Reproducibility: Stratification uses a fixed random seed (42) with proportional allocation by culture size. Per-sample audit outcomes (pass/fail/corrected) are logged in audit_log_p3.csv (supplementary), and full reviewer assignments are documented in Appendix E.
Table 6 reports quality metrics with Wilson confidence intervals: 100% bilingual completeness, 0% duplicates, 100% image availability (100-sample audit), and 100% dimension coverage ($\geq$70% threshold). To address sample-size imbalance concerns, we provide a balanced subset (N=48 per culture $\times$ 7, total 336 pairs). Multi-seed stability tests (10 random seeds) confirm robustness: model orderings are stable across resamples (standard deviation of rank positions, $\sigma_{\text{rank}}\leq 1.5$). Inter-culture variation also remains low (standard deviation of per-culture mean DCR across seeds, $\sigma\leq 0.05$). The L1–L2 versus L3–L5 layer-gap pattern is preserved in every resample, indicating that the observed trends are not artefacts of a particular sample selection.
Metric Val 95% CI
Bilingual complete 100% [99.9%, 100%]
Duplicate rate 0% [0%, 0.1%]
Image available 100% [96.3%, 100%]
Dim $\geq$70% 100% [99.9%, 100%]
Table 6: Quality metrics with Wilson CI. 100-sample audit.
Vulca-Bench supports multiple research applications. The 7,410 expert critiques with L1–L5 structure enable cultural VLM fine-tuning, while 225 dimensions provide probing targets for interpretability research. The balanced subset (N=336 for pilot, 48/culture $\times$ 7; N=384 available for full evaluation with 8 cultures) enables equal-weighted cultural fairness probing, and bilingual critiques support cultural knowledge retrieval and RAG augmentation. [^1] Hermitage excluded from pilot due to late inclusion; full 8-culture evaluation (N=384) available in released dataset.
6 Pilot Evaluation
To demonstrate Vulca-Bench’s utility as a cultural VLM benchmark, we evaluate 5 representative VLMs on a stratified sample. This pilot evaluation establishes dataset utility and diagnostic capability only; it does not constitute a definitive evaluation framework. We use Dimension Coverage Rate (DCR) as a dataset-level diagnostic check, verifying that VLM outputs engage with culture-specific concepts, rather than as a leaderboard metric for model ranking.
6.1 Setup
We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash (temperature=0.7, max_tokens=2048; full configuration in supplementary). We use Dimension Coverage Rate (DCR) as a diagnostic to test whether model-generated critiques engage with culture-specific concepts. For a critique $c$ from culture $k$:
$\text{DCR}(c,k)=\frac{|\mathcal{D}_{k}^{c}|}{|\mathcal{D}_{k}|}$ (1)
where $\mathcal{D}_{k}$ is the culture-specific dimension set (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$) and $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ denotes the subset of dimensions detected via keyword matching. Each record’s covered_dimensions field (Section 3) provides the gold set, namely dimension IDs explicitly labeled by annotators during critique creation. All models generate English critiques; DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
DCR is intended as a coarse diagnostic rather than a leaderboard metric. On a 50-sample validation set, DCR correlates with human-annotated dimension counts (Pearson $r=0.82$). Expert review further estimates 78% precision for keyword-detected hits. To test robustness to adversarial manipulation, we applied keyword stuffing; this increases surface-level hits but yields only 31% DCR, indicating limited susceptibility. For semantic validation, we report agreement rates between keyword detections and two independent checks at the sample level: (i) embedding-based similarity (cosine $\geq 0.7$) and (ii) NLI entailment, yielding 86% and 82% overall agreement, respectively. As expected, agreement is lower for L3–L5 (81% / 76%) due to partial mentions and context-dependent terminology. Full protocol details (thresholds, dictionaries, and length-controlled analyses) appear in Appendix D.
6.2 Dimension Coverage Results
Model L1–L2 L3–L5 $\Delta_{L}$ DCR
Gemini-2.5-Pro 89.2 58.1 31.1 72.4
GPT-4o 87.1 46.8 40.3 65.3
Qwen3-VL-235B 85.6 54.3 31.3 68.7
Claude-Sonnet-4.5 84.3 48.2 36.1 64.8
GLM-4V-Flash 78.4 40.7 37.7 58.2
Table 7: Layer-Gap Diagnostic results. $\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit. All models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5). These values should be interpreted as a reference rather than state-of-the-art results because DCR is a coarse proxy and the pilot uses a limited prompt configuration. N=336 balanced-pilot subset (7 cultures); bootstrap 95% CI half-width $\approx$ 4.8 pp.
All models exhibit a 25–40 percentage-point layer gap ($\Delta_{\text{L}}=\text{DCR}_{\text{L1--L2}}-\text{DCR}_{\text{L3--L5}}$), confirming Vulca-Bench captures cultural understanding beyond visual perception. The 14-point DCR spread (72.4%–58.2%) demonstrates discriminative power across model capabilities. Critically, model orderings on the balanced-pilot subset (N=336, 7 cultures) correlate strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C). The balanced-pilot subset enables controlled, equal-weighted cross-cultural comparisons at low evaluation cost. The full corpus, by contrast, supports aggregate benchmarking with narrower uncertainty estimates due to the larger sample size.
6.3 Implications
These results confirm our hypothesis that cultural understanding is fundamentally hierarchical. The consistent 25–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability tier from visual perception. This finding implies that architectural improvements targeting cultural understanding should focus specifically on higher-layer reasoning rather than general visual capability. We release and validate DCR (78% precision; resistant to adversarial manipulation) as a reproducible, dataset-level diagnostic check signal for dimension coverage.
6.4 Error Analysis: VLM Failures on L3–L5
Beyond aggregate metrics, Vulca-Bench enables fine-grained error diagnosis. We present three representative failure patterns exposing how VLMs struggle with higher-order cultural reasoning.
We identify three recurring error patterns (detailed examples in Appendix H): (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c. vanitas to 16th-c. pronkstilleven); (3) Cultural conflation—confusing distinct traditions despite clear stylistic markers (e.g., Persian vs. Mughal miniatures). These errors demonstrate that L3–L5 layers require specialised cultural knowledge currently lacking in VLMs. Vulca-Bench’s 195 dimensions provide diagnostic precision to categorise such failures systematically.
6.5 Few-Shot Learning with Expert Critiques
We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs). We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars (Table 8).
Contrary to expectations, few-shot prompting yields decreased performance across all tested VLMs. DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$). GPT-4o follows with $15.5\%$ decline ($p<0.001$). Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
We hypothesise three contributing factors: (1) attention dilution—longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting—exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity—exposure to expert critique structures may constrain flexibility. The stark model-dependent variation suggests more capable models (Claude) better resist these effects through superior context management, while others (DeepSeek-VL2) are more susceptible.
[…]
Model N 0-shot 1-shot 3-shot $\Delta$ $p$
DeepSeek-VL2 362 0.985 0.819 0.578 -41.3% <<0.001
GPT-4o 338 0.998 0.967 0.843 -15.5% <<0.001
Claude-Opus 183 1.000 0.998 0.986 -1.4% 0.074
Claude-Sonnet 145 1.000 0.992 0.985 -1.5% 0.124
Table 8: Few-shot prompting performance (DCR) across 4 VLMs on Chinese and Western art (NN=1,028). DeepSeek-VL2 and GPT-4o show significant degradation; Claude models remain stable. $\Delta$: relative change from 0-shot to 3-shot. Significance: *** $p<0.001$.
7 Release and Licensing
The complete dataset is available under CC BY 4.0 in the supplementary materials. The release package includes: metadata JSON (7,410 pairs with bilingual critiques and dimension annotations), all images (optimised to $\leq$3.75MB), 225 dimension definitions, synonym dictionaries, evaluation scripts, and annotation guidelines. We release the full evaluation package and provide cached model outputs to support reproducibility. Using cached responses, the pilot experiments can be replicated at under USD 20 in API costs at the time of writing (see reproduce_pilot.py for the exact configuration). All artworks are from public museum collections (public domain or CC0); each record includes source_institution and accession_number for provenance (see Appendix I for culture distribution visualization).
8 Conclusion
We present Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, operationalised through a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions. The dataset addresses critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
Vulca-Bench enables systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception. The complete dataset (images, metadata, critiques, dimension definitions, and evaluation/annotation tools) is available under CC BY 4.0 in the supplementary materials. We emphasise that DCR serves as a coarse-grained diagnostic rather than a leaderboard metric.
Vulca-Bench enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset. The dataset complements, not replaces, human cultural expertise.
Limitations
Our work has several limitations. First, regarding corpus distribution, Chinese and Western traditions dominate (82%), reflecting real-world museum digitisation availability and expert accessibility rather than methodological bias. This natural distribution may cause higher variance in evaluation results for minority cultures, so cross-cultural fairness analyses should default to the balanced-pilot subset (N=336, 7 cultures) or full balanced subset (N=384, 8 cultures) with per-culture CI reporting (see Appendix B).
Second, L5 (philosophical aesthetics) inherently involves interpretive judgment, with higher reviewer correction rates at L5 (3.8%) compared to L1–L2 (0.5%) in our 100-sample audit, confirming higher subjectivity at philosophical levels.
Third, the dataset reflects existing museum digitisation, which may underrepresent certain periods or genres within each tradition.
Fourth, our Cultural Symmetry evidence (layer difficulty $r=0.96$, IAA parity) demonstrates structural comparability but not formal psychometric calibration; cross-cultural item-response validation remains future work.
Fifth, our Chinese–English bilingual design, while enabling cross-cultural accessibility and preserving key Chinese aesthetic terminology, may introduce translation loss for non-Chinese traditions. Japanese concepts (wabi-sabi, mono no aware), Korean aesthetics (jeong, heung), Islamic calligraphic principles, and Indian philosophical frameworks (rasa, bhava) possess native terminologies that English romanisation only partially captures. This limitation reflects expert availability constraints rather than linguistic preference; extending to native-language critiques (Japanese, Korean, Arabic, Hindi/Sanskrit) is an important future direction. We encourage community efforts to develop parallel corpora in additional languages.
Finally, DCR serves as a coarse diagnostic rather than a precision instrument, designed for dataset utility validation rather than definitive model ranking.
Ethical Considerations.
All artworks are sourced from public museum collections with open-access policies (public domain or CC0), and the complete dataset is available under CC BY 4.0 in the supplementary materials, with no copyrighted contemporary art included. Expert annotators from each cultural tradition ensure respectful and accurate representation; we acknowledge that computational cultural analysis cannot replace human expertise, and our goal is to assist, not supplant, cultural custodians.
Regarding generative misuse risk, expert critique data could potentially be used to train VLMs that generate fake “expert” art commentary, risking misinformation in cultural education. We mitigate this through source attribution (all dataset JSON records include provenance metadata linking critiques to Vulca-Bench), detection research (we encourage using our bilingual critiques as positive examples for training AI-generated content detectors), and usage terms (CC BY 4.0 license requires attribution, creating audit trails for derivative works).
Appendix A Dimension Definitions
Full dimension definitions for each culture are provided in the supplementary materials. Example dimensions:
Chinese (30 dimensions).
  • CN_L1_D1: Color Palette (设色)
  • CN_L3_D2: Symbolic Motifs (象征寓意)
  • CN_L5_D1: Spirit Resonance (气韵生动)
Western (25 dimensions).
  • WE_L1_D1: Composition and Balance
  • WE_L3_D2: Iconographic Analysis
  • WE_L5_D1: Art Historical Significance
Hermitage (30 dimensions).
  • WS_L1_D1: Color Palette (色彩)
  • WS_L3_D1: Religious Iconography (宗教图像)
  • WS_L5_D6: Cultural Heritage (文化遗产)
Appendix B Dataset Variants
Table 9 clarifies the relationship between dataset variants. Key boundary: Full corpus and Balanced include all 8 cultures (with Hermitage); Eval uses 7 cultures (Hermitage excluded); Gold/Human subsets use 6 cultures (Mural and Hermitage excluded due to insufficient gold annotations for reliable Tier III calibration).
Variant Pairs C D Usage
Full 7,410 8 225 Complete corpus
Eval 7,214 7 195 VLM evaluation (excl. Hermitage)
Gold 294 6 165 Expert references
Human 450 6 Tier III calibration
Balanced 384 8 225 Fairness analysis (48/culture)
Balanced-Pilot 336 7 195 Pilot evaluation (excl. Hermitage)
Table 9: Dataset variants. C=cultures, D=dimensions. 8-culture: Full, Balanced (include Hermitage). 7-culture: Eval, Balanced-Pilot (exclude Hermitage). 6-culture: Gold, Human (exclude Mural, Hermitage due to insufficient gold annotations). Hash b8a34e5f* verifies single source.
Appendix C Full-Corpus Ordering Consistency
Table 10 validates that model orderings on the balanced-pilot subset (N=336, 7 cultures) generalise to the full corpus (N=7,410, 8 cultures). Spearman $\rho=0.94$ confirms high rank correlation; bootstrap 95% CI [0.87, 0.98] excludes chance agreement.
Model Bal. Full B-Rk F-Rk
[…]
Appendix D DCR Validation Details
Length-controlled analysis.
To ensure DCR measures semantic coverage rather than critique length, we performed length-normalised analysis: correlation between critique length and DCR is $r=0.23$, indicating length contributes minimally to scores. Models with similar output lengths (GPT-4o: 312 words avg, Claude: 298 words avg) show 6-point DCR differences, confirming semantic content drives variation.
Synonym dictionary.
We maintain culture-specific synonym dictionaries (e.g., “chiaroscuro” $\approx$ “light-dark contrast”, “impasto” $\approx$ “thick paint application”). Dictionary v2.1 (current) includes 847 term mappings across 7 cultures, validated by domain experts. Version history and updates are tracked in supplementary materials.
Semantic alignment.
Embedding-based validation (text-embedding-3-small) confirms keyword hits correlate with semantic similarity: 86% overall agreement between DCR keyword detection and embedding cosine similarity $\geq$0.7. L3–L5 shows lower alignment (81%) due to partial mentions and context-dependent terminology.
Appendix E Annotation Guidelines
Annotators follow a standardised protocol:
  1. 1.
    View artwork image at full resolution
  2. 2.
    Identify artist, period, and cultural context
  3. 3.
    Write critique covering all 5 layers
  4. 4.
    Ensure $\geq$70% dimension coverage
  5. 5.
    Translate to English preserving cultural terms
Appendix F Sample Critiques
We present representative bilingual critiques demonstrating L1–L5 coverage.
Chinese Example: Dunhuang Mural Hand Gestures.
Western Wei Dynasty, Mogao Cave 249
Chinese (excerpt): 此组莫高窟249窟西魏壁画手势白描稿,聚焦于佛教造像的手印细节刻画.线描技法运用高古游丝描,线条细劲均匀,转折圆润,富于弹性韵律(L2).手势造型遵循佛教图像学规范,如合十代表敬礼,施无畏印表示护佑(L3).西魏时期敦煌壁画受北朝遗风影响,线条风格较为古拙刚健(L4).
English (excerpt): This set of Western Wei mural hand-gesture drafts from Mogao Cave 249 focuses on Buddhist mudras. The line technique employs gaogu yousi miao (ancient silk-thread strokes), with fine, even lines and rounded turns (L2). Hand gestures follow Buddhist iconographic conventions: anjali mudra represents reverence, abhaya mudra signifies protection (L3). Western Wei Dunhuang murals show Northern Dynasties influence with archaic, vigorous line quality (L4).
Dimensions covered: CN_L1_D3, CN_L2_D1, CN_L2_D5, CN_L3_D1, CN_L4_D1
Western Example: Impressionist Landscape.
“Paysage de Saint-Cheron” by J.-B. Armand Guillaumin
Chinese (excerpt): 吉约曼是法国印象派重要成员,与塞尚,毕沙罗交往密切.此作展现其标志性的浓烈色彩与奔放笔触,色彩饱和度高于其他印象派画家.光影处理大胆,明暗对比强烈,营造戏剧性视觉效果(L1-L2).其风格介于印象派与后印象派之间,对色彩表现力的强调预示了野兽派等后续流派的发展(L5).
English (excerpt): Guillaumin was a key member of French Impressionism, closely associated with Cézanne and Pissarro. This work demonstrates his signature intense colors and bold brushwork, with higher saturation than other Impressionists. Light handling is dramatic with strong contrast (L1–L2). His style bridges Impressionism and Post-Impressionism, anticipating Fauvism through emphasis on color expressivity (L5).
Dimensions covered: WE_L1_D1, WE_L2_D3, WE_L4_D2, WE_L5_D1
Japanese Example: Ukiyo-e Warrior Print.
“Three Warriors” by Utagawa Kuniyoshi, Edo Period
Chinese (excerpt): 歌川国芳的这幅浮世绘木版画展现三位武士的动态构图,体现江户时代武者绘典型特征.画面采用对角线构图,营造强烈戏剧张力(L1).色彩以蓝,黑,红为主,金黄点缀盔甲细节(L2).画面上方题跋与下方视觉中心形成虚实对比,体现"间"(ma)的美学理念(L5).
English (excerpt): This ukiyo-e woodblock print by Utagawa Kuniyoshi depicts three warriors in dynamic composition, embodying Edo period musha-e characteristics. Diagonal composition creates dramatic tension (L1). Color palette employs blues, blacks, and reds with golden accents on armor (L2). Upper inscriptions contrast with lower visual center, demonstrating the aesthetic concept of ma (negative space) (L5).
Dimensions covered: JP_L1_D2, JP_L2_D1, JP_L3_D4, JP_L5_D2
Appendix G Open-Source Embedding Replication
To ensure reproducibility without proprietary API dependencies, we replicated the semantic alignment analysis (Section 6) using the open-source BGE-large-en-v1.5 model. [^2] Xiao et al., 2023. C-Pack: Packaged Resources To Advance General Chinese Embedding. https://arxiv.org/abs/2309.07597
Layer OpenAI BGE $\Delta$
L1–L2 94% 92% –2%
L3–L5 81% 79% –2%
Overall 86% 84% –2%
Table 11: Semantic alignment comparison: OpenAI text-embedding-3-small vs BGE-large-en-v1.5 (open-source). Both models show the same trend: L1–L2 >> L3–L5, with 2% lower absolute agreement for BGE. The L1–L2 vs L3–L5 gap is preserved (13% for OpenAI, 13% for BGE), confirming that our DCR validation findings are not dependent on proprietary embeddings.
Reproducibility: BGE-large-en-v1.5 is available on Hugging Face (BAAI/bge-large-en-v1.5) under MIT license. We release the replication script (eval/semantic_align_bge.py) with the dataset.
Appendix H Error Pattern Examples
Surface-level terminology.
GPT-4o analysing a Song Dynasty landscape scroll produces: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations. Expert correction identifies three visual strategies: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs. dian), and strategic negative space (liubai).
Historical anachronism.
Claude-Sonnet-4.5 on a 16th-century Dutch market scene: “This vanitas painting conveys Protestant anxieties about mortality… characteristic of 17th-century Dutch Golden Age memento mori.” The VLM applies 17th-century conventions to a 1560s pronkstilleven emphasizing Renaissance prosperity rather than later Calvinist meditation.
Cultural conflation.
Gemini-2.5-Pro identifies a Safavid Persian miniature as “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
Appendix I Culture Distribution Visualization
Figure 3: Dataset Distribution. Culture distribution across 8 traditions (7,410 pairs). Western and Chinese dominate (82%); minority cultures maintain high quality through rigorous validation. Complete dataset is available under CC BY 4.0 in supplementary materials.

### v2

@@
 Our pilot results indicate that higher-layer reasoning (L3–L5) is consistently more challenging than visual and technical analysis (L1–L2).
-The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 in the supplementary materials.
+The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 on GitHub [^1] https://github.com/yha9806/VULCA-Bench and Hugging Face [^2] https://huggingface.co/datasets/harryHURRY/vulca-bench.
 Vulca-Bench: A Multicultural Vision-Language Benchmark
 for Evaluating Cultural Understanding
-Haorui Yu and Ramon Ruiz-Dolz and Diji Yang and Hang He and Fengrui Zhang and Qiufeng Yi
+Haorui Yu1, Ramon Ruiz-Dolz2, and Diji Yang3, and Hang He4, and Fengrui Zhang5, and Qiufeng Yi6 1DJCAD, University of Dundee, United Kingdom 2Faculty of Science, Engineering and Business, University of Dundee, United Kingdom 3 University of California, Santa Cruz, USA 4East China Normal University, China 5Nanjing University, China 6School of Computer Science, University of Birmingham, United Kingdom
 1 Introduction
-Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks, including 90%+ on object detection Li et al.
-(2023) and 77.2% on VQAv2 Goyal et al.
-(2017).
-However, these benchmarks predominantly assess L1–L2 visual perception, which involves recognising objects, describing scenes, and answering factual questions.
+Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks.
+For instance, on object hallucination benchmarks like POPE Li et al.
+(2023) where leading models such as InstructBLIP reach 88.7% accuracy, and on visual question answering such as VQAv2 Goyal et al.
+(2017) where best models exceed 75% accuracy.
+However, these benchmarks predominantly assess visual perception, which involves recognising objects, describing scenes, and answering factual questions.
 They fail to evaluate cultural understanding, namely the capacity to interpret symbolic meanings, appreciate aesthetic traditions, and engage with philosophical concepts embedded in visual content.
-Recent studies confirm this gap: GPT-4o achieves only 54.1% accuracy on Chinese traditional painting Zeng et al.
+Recent studies confirm this gap: GPT-4o achieves only 54.1% overall accuracy on CII-Bench, with even lower performance (51.8%) on Chinese Traditional Culture content Zhang et al.
 (2025), and cross-cultural evaluations reveal significant Western bias in VLM performance Liu et al.
 (2025).
-We define cultural understanding using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics).
-Higher layers require progressively deeper cultural knowledge (Section 3 provides full definitions).
-Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3) Zhang (2024), the artist’s lineage within the “Four Gentlemen” tradition (L4), or the aesthetic principle of 气韵生动 (qiyun shengdong, “spirit resonance”) that defines Chinese painting philosophy (L5) Jing (2023).
+We define cultural understanding by using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics).
+Higher layers require progressively deeper cultural knowledge (see Section 3 for complete definitions).
+Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3), the artist’s lineage within the “Four Gentlemen” (四君子) tradition (L4), or aesthetic principles such as 气韵生动 (qiyun shengdong, “spirit resonance”) and 意境 (yijing, “artistic conception”) Jing (2023) that define Chinese painting philosophy (L5).
 This example illustrates that cultural understanding is not a single capability but a progression across interpretive levels, a progression that existing benchmarks fail to measure.
-Recent cultural AI datasets (CulturalBench Chiu et al.
+Recent cultural AI datasets (e.g., CulturalBench Chiu et al.
 (2025), CultureAtlas Fung et al.
@@
 Initial work on Chinese painting critique Yu et al.
-(2025a) introduced quantitative VLM evaluation using 163 expert commentaries, revealing significant VLM-expert divergence.
+(2025a) introduced a quantitative VLM evaluation method using 163 expert commentaries, revealing significant VLM-expert divergence.
 Complementary probing Yu et al.
-(2025b) found VLMs rely on “symbolic shortcuts,” excelling on Western festivals but failing on underrepresented traditions.
+(2025b) found VLMs rely on “symbolic shortcuts”, excelling on Western festivals but failing on under-represented traditions.
 However, these studies focused on single cultures; systematic cross-cultural evaluation remains absent.
-Existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
-We introduce Vulca-Bench (Vision-Understanding-Language-Culture Assessment), a comprehensive multicultural art critique benchmark addressing these gaps.
-Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and 98% cultural-fact accuracy, operationalised through our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size).
+Furthermore, existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
+This paper introduces Vulca-Bench (Vision-Understanding and Language-based Cultural Assessment), a comprehensive multicultural art critique benchmark addressing these gaps.
+Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and expert-validated factual accuracy, annotated using our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size).
 All critiques meet expert-quality standards (Chinese $\geq$150 characters, English $\geq$100 words) with cultural specialist validation.
-The complete dataset is available under CC BY 4.0 in the supplementary materials.
 Vulca-Bench enables systematic evaluation of VLMs’ hierarchical cultural understanding, identification of L1–L5 performance gaps, equal-weighted cultural fairness probing on the balanced-pilot subset (N=336, 7 cultures), and development of culturally grounded VLM architectures.
 Figure 1 showcases representative artworks from each tradition.
-This paper makes three contributions.
+This paper therefore makes three contributions.
 (1) We present Vulca-Bench, a multicultural art-critique benchmark enabling evaluation of cultural interpretation beyond surface-level perception, comprising 7,410 image–critique pairs across eight traditions with 225 expert-defined dimensions.
@@
 Standard VLM benchmarks focus on L1 perception: MME Fu et al.
-(2023) and SEED-Bench-2 Li et al.
-(2024) achieve 90%+ on object detection but lack cultural evaluation; POPE Li et al.
-(2023) tests object existence without symbolic significance; and specialised benchmarks such as MathVista Lu et al.
+(2025) evaluates 14 subtasks spanning perception and cognition, SEED-Bench-2 Li et al.
+(2024) covers 27 dimensions with 24K multiple-choice questions, and POPE Li et al.
+(2023) probes object hallucination through binary yes/no questions.
+However, none address cultural understanding in-depth.
+Specialised benchmarks such as MathVista Lu et al.
 (2024), ChartQA Masry et al.
 (2022), and DocVQA Mathew et al.
-(2021) evaluate domain reasoning but not cultural understanding.
+(2021) evaluate domain reasoning but not cultural interpretation.
 Cultural AI datasets have begun to address this gap, yet remain limited.
@@
 (2024) adopt QA formats rather than generative critique; GIMMICK Schneider et al.
-(2025) reveals Western biases across 144 countries but employs recognition tasks; WuMKG Wan et al.
-(2024) provides a Chinese painting knowledge graph (104K entities) for retrieval, not generative evaluation; and CulTi Yuan et al.
+(2025) reveals Western biases across 144 countries employing recognition tasks; WuMKG Wan et al.
+(2024) provides a Chinese painting knowledge graph (104K entities) useful for information retrieval rather than generative evaluation; and CulTi Yuan et al.
 (2025) offers image-text pairs for heritage retrieval without hierarchical cultural dimensions.
 Art and aesthetics datasets similarly fall short.
-WikiArt Saleh and Elgammal (2015) (77K paintings, 27 styles) and OmniArt Strezoski and Worring (2018) (429K artworks) enable style classification but lack expert critiques or cultural dimension annotations; ArtEmis Achlioptas et al.
+WikiArt Saleh and Elgammal (2016) ($\sim$80K paintings, 27 styles) and OmniArt Strezoski and Worring (2018) (>>1M artworks) enable style classification but lack expert critiques or cultural dimension annotations; ArtEmis Achlioptas et al.
 (2021) focuses on affective responses; and Zhang et al.
-Zhang (2024) model Chinese painting critique via Xie He’s Six Principles at L1–L2 only.
+(2024) survey computational approaches for traditional Chinese painting based on Xie He’s Six Principles, focusing primarily on visual and technical analysis.
 Prior VLM art critique work Yu et al.
@@
 Vulca-Bench addresses all three gaps through cross-cultural extension to eight traditions.
-Unlike QA-based cultural benchmarks that test factual recall, or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5).
-Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (layer-wise drop analysis, dimension coverage scoring) unavailable in existing datasets.
+Unlike QA-based cultural benchmarks that test factual recall or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5).
+Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (e.g., layer-wise analysis or dimension coverage scoring) unavailable in existing datasets.
 3 Dataset Construction
@@
 This design reduces Western-centric methodological bias while acknowledging that museum availability and expert access naturally produce imbalanced sample sizes.
-Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage).
-We recommend using balanced variants for per-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
+Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage, whose European court art overlaps with Western).
+We recommend using balanced variants for cross-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
 We operationalise Cultural Symmetry through three pillars.
@@
 We therefore set this threshold empirically, based on the observation that lower coverage correlated strongly with missing philosophical or historical analysis.
-Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates, distributed proportionally across cultures).
+Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates distributed proportionally across cultures).
 Beyond structural parity, we validate measurement alignment using three criteria.
-First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures; the resulting five-point layer profiles are highly similar (mean pairwise Pearson correlation across cultures $r=0.96$).
+First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures.
+All cultures show the same pattern: L1 easiest, L5 hardest ($r=0.96$ cross-culture correlation).
 Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.79$ for L1–L2 and $0.71$ for L3–L5, with cross-culture standard deviation below 0.04.
@@
 We operationalise cultural understanding through our five-layer framework reflecting increasing interpretive complexity.
-L1–L2 require visual observation; L3–L5 demand progressively deeper cultural knowledge that VLMs typically lack.
+While L1–L2 require visual observation, L3–L5 demand progressively deeper cultural and contextual knowledge.
 Layer Description
@@
 Each culture adapts L1–L5 to its indigenous aesthetic vocabulary, yielding 225 dimensions total.
-Dimension counts (25–30) reflect tradition complexity: Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境); Western 25 dimensions emphasise formal analysis (chiaroscuro, perspective).
-All critiques must cover $\geq$70% of culture-specific dimensions, ensuring substantive L3–L5 engagement.
+Dimension counts (25–30) reflect tradition complexity.
+For example, Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境).
+On the other hand, Western cultures cover 25 dimensions that emphasise more formal analysis (chiaroscuro, perspective).
+In Vulca-Bench, all critiques must cover at least 70% of their respective culture-specific dimensions, ensuring substantive L3–L5 engagement.
 Culture Dims Pairs Key Concepts
-Western 25 4,041 Chiaroscuro, linear perspective, impasto
+Western 25 4,041 Chiaroscuro, perspective, impasto
 Chinese 30 2,042 气韵 (qiyun), 笔墨 (bimo), 意境 (yijing)
-Japanese 27 401 Wabi-sabi, yugen, ma (negative space)
-Islamic 28 205 Geometric patterns, arabesques, Persian miniature
-Mural 30 201 Buddhist art, Dunhuang murals, cave painting
-Hermitage 30 196 Russian/European court art, iconography
-Indian 30 173 Rasa theory (9 emotions), shringara, bhakti
-Korean 25 151 Literati aesthetics, muninhwa tradition
+Japanese 27 401 Wabi-sabi, yugen, ma
+Islamic 28 205 Geometric patterns, arabesques
+Mural 30 201 Buddhist art, Dunhuang murals
+Hermitage 30 196 Russian/European court art
+Indian 30 173 Rasa theory, shringara, bhakti
+Korean 25 151 Literati aesthetics, muninhwa
 Total 225 7,410 8 cultures, Cultural Symmetry
@@
 Chinese artworks derive from the Palace Museum, Shanghai Museum, and Metropolitan Museum; Western works from the Metropolitan Museum, Louvre, and Hermitage Museum; Japanese from Tokyo and Kyoto National Museums; Korean from the National Museum of Korea; Islamic from the Museum of Islamic Art Doha and Metropolitan Museum; Indian from the National Museum Delhi and British Museum; and Hermitage works from the State Hermitage Museum collection (Russian and European court art, religious iconography).
-All images are released under CC BY 4.0 in the supplementary materials.
+All images are released under CC BY 4.0 at https://github.com/yha9806/VULCA-Bench and HuggingFace: https://huggingface.co/datasets/harryHURRY/vulca-bench.
 3.4 Bilingual Critique Annotation
@@
 Third, expert availability: our annotation team (Table 3) comprises bilingual specialists with deep cultural expertise, enabling controlled drafting, translation, and review within a single protocol.
-This design reflects practical constraints rather than linguistic preference; extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in Limitations.
+This design reflects practical constraints rather than linguistic preference, thus extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in the Limitations section.
 Culture Ann. Rev. Background
@@
 A reviewer validates each critique on dimension coverage ($\geq$70%), cultural accuracy, and bilingual consistency.
-Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources; unresolved cases (<<2%) escalate to senior expert panel.
-Full annotation guidelines appear in Appendix E; representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
+Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources.
+The complete scheme of our annotation guidelines can be found in Appendix E and representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
 We measure inter-annotator agreement on a 100-sample subset (Table 4).
@@
 Korean follows a modified protocol due to the scarcity of qualified Korean art-history specialists with dual fluency.
-Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with 100% expert review by a senior scholar.
-The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA; we report Korean separately for methodological transparency.
+Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with a 100% expert review by a senior scholar.
+The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA.
 Culture N L1–L2 $\kappa$ L3–L5 $\kappa$ ICC
@@
 Korean uses a different protocol: 1 primary annotator (PhD Korean studies) + 100% expert review.
-The reported $\kappa$ is annotator–reviewer agreement, not independent dual-annotation IAA; we report it separately for quality transparency but do not include it in pooled statistics.
-3.5 Quality Assurance
-Multi-stage validation ensures data quality.
-Each record contains an explicit dimension checklist through its covered_dimensions field, listing dimension IDs (e.g., ["CN_L1_D1", "CN_L3_D2", ...]) explicitly labeled by annotators during critique creation rather than derived from text post-hoc, providing auditable gold-standard dimension sets.
-Automated checks enforce the $\geq$70% dimension coverage threshold (21/30 for CN, 18/25 for WE, etc.) based on covered_dimensions length.
-We removed all exact duplicates after deep audit.
-Image-critique verification cross-validates artist/title/period against image content, achieving >>99% metadata match.
-All pairs have both Chinese ($\geq$150 characters) and English critiques, ensuring bilingual completeness.
-Finally, all images are compressed to $\leq$3.75MB for VLM API compatibility.
 4 Dataset Statistics
 Vulca-Bench comprises 7,410 matched image-critique pairs drawn from 11,964 total images across 8 cultural traditions, annotated with 225 culture-specific dimensions.
-Table 5 summarises key metrics: average Chinese critique length is 450 characters (English: 280 words), with 100% bilingual completeness, zero duplicates, >99% metadata verification, and 98% cultural-fact accuracy.
+Table 5 summarises key metrics.
+The average Chinese critique length is 450 characters, while English critiques have an average of 280 words.
+Bilingual coverage was complete with all samples having EN-CN critiques, and all duplicates were excluded.
 All pairs meet the $\geq$70% dimension coverage threshold.
@@
 […]
-Figure 2 reveals layer-wise coverage patterns: L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), while L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding.
+Figure 2 reveals layer-wise coverage patterns.
+While L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding.
 Chinese art shows highest L5 coverage (78%), consistent with its rich philosophical aesthetics tradition (qiyun, yijing).
-Importantly, each culture’s critiques are authored by native expert annotators from that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
-L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapanese
-Figure 2: Layer-wise dimension coverage. L1–L2 achieve $\geq$94%; L3–L5 decline to 72–89%, reflecting interpretive complexity.
-5 Data Quality Validation
-We validate Vulca-Bench through a three-phase audit protocol.
-Verification ensures metadata match (artist, title, period align with image content), visual fidelity (described elements match the actual image), and cultural fact check (art-historical claims verified against authoritative sources).
-Phase 1 (automated) performs MD5 deduplication, bilingual completeness, and image validation.
-Phase 2 (stratified) constitutes the 450 human-scored subset.
-KR: 107 (full set), JP/IS/IN: 50 each, CN/WE: 100 each.
-Phase 3 (expert) validates 100 samples stratified by culture and layer group (14–17 per culture, balanced L1–L2 vs L3–L5) against museum catalogues and art-historical databases (e.g., Grove Art Online, Oxford Art Online).
-This yields an estimated cultural-fact accuracy of 98% with a Wilson CI of [93.0%, 99.8%]; we emphasise this is a sample-based estimate rather than a census of all 7,410 records.
-Reproducibility: Stratification uses a fixed random seed (42) with proportional allocation by culture size.
-Per-sample audit outcomes (pass/fail/corrected) are logged in audit_log_p3.csv (supplementary), and full reviewer assignments are documented in Appendix E.
-Table 6 reports quality metrics with Wilson confidence intervals: 100% bilingual completeness, 0% duplicates, 100% image availability (100-sample audit), and 100% dimension coverage ($\geq$70% threshold).
-To address sample-size imbalance concerns, we provide a balanced subset (N=48 per culture $\times$ 7, total 336 pairs).
-Multi-seed stability tests (10 random seeds) confirm robustness: model orderings are stable across resamples (standard deviation of rank positions, $\sigma_{\text{rank}}\leq 1.5$).
-Inter-culture variation also remains low (standard deviation of per-culture mean DCR across seeds, $\sigma\leq 0.05$).
-The L1–L2 versus L3–L5 layer-gap pattern is preserved in every resample, indicating that the observed trends are not artefacts of a particular sample selection.
-Metric Val 95% CI
-Bilingual complete 100% [99.9%, 100%]
-Duplicate rate 0% [0%, 0.1%]
-Image available 100% [96.3%, 100%]
-Dim $\geq$70% 100% [99.9%, 100%]
-Table 6: Quality metrics with Wilson CI. 100-sample audit.
-Vulca-Bench supports multiple research applications.
-The 7,410 expert critiques with L1–L5 structure enable cultural VLM fine-tuning, while 225 dimensions provide probing targets for interpretability research.
-The balanced subset (N=336 for pilot, 48/culture $\times$ 7; N=384 available for full evaluation with 8 cultures) enables equal-weighted cultural fairness probing, and bilingual critiques support cultural knowledge retrieval and RAG augmentation.
-[^1] Hermitage excluded from pilot due to late inclusion; full 8-culture evaluation (N=384) available in released dataset.
+Importantly, each culture’s critiques are authored by annotators with specialized expertise in that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
+L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapaneseKoreanIslamicIndianMuralHermitage
+Figure 2: Layer-wise dimension coverage across all right cultures. L1–L2 achieve $\geq$93%; L3–L5 decline to 70–89%, reflecting interpretive complexity.
 5 Pilot Evaluation
+Gemini-2.5-ProQwen3-VL-235BGPT-4oClaude-Sonnet-4.5GLM-4V-Flash02020404060608080100100Coverage (%)(a) Layer-wise CoverageL1–L2 (Visual/Technical)L3–L5 (Cultural/Philosophical)DeepSeek-VL2GPT-4oClaude-OpusClaude-Sonnet05050100100Coverage (%)(b) Few-shot Prompting Results0-shot1-shot3-shot
+Figure 3: VLM diagnosis reveals systematic cultural understanding gaps.
+(a) Layer-wise coverage: all models exhibit 31–40 percentage-point gaps between L1–L2 and L3–L5, confirming cultural/philosophical layers are harder than visual/technical analysis.
+(b) Few-shot prompting results: DeepSeek-VL2 and GPT-4o show significant degradation from 0-shot to 3-shot; Claude models remain robust.
 To demonstrate Vulca-Bench’s utility as a cultural VLM benchmark, we evaluate 5 representative VLMs on a stratified sample.
-This pilot evaluation establishes dataset utility and diagnostic capability only; it does not constitute a definitive evaluation framework.
-We use Dimension Coverage Rate (DCR) as a dataset-level diagnostic check, verifying that VLM outputs engage with culture-specific concepts, rather than as a leaderboard metric for model ranking.
+This pilot evaluation is aimed at showcasing dataset utility and does not represent a definitive benchmark to compare against.
+For that purpose, we use Dimension Coverage Rate (DCR) to verify that VLM outputs engage with culture-specific concepts and to highlight the usefulness of Vulca-Bench, specifically focusing on cultural coverage rather than on a complete assessment of the VLM-generated critiques.
 5.1 Setup
-We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash (temperature=0.7, max_tokens=2048; full configuration in supplementary).
-We use Dimension Coverage Rate (DCR) as a diagnostic to test whether model-generated critiques engage with culture-specific concepts.
+We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash.
+Temperature was set to 0.7 and max_tokens to 2048.
+We then use DCR as a measure to test whether model-generated critiques engage with culture-specific concepts.
 For a critique $c$ from culture $k$:
@@
 where $\mathcal{D}_{k}$ is the culture-specific dimension set (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$) and $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ denotes the subset of dimensions detected via keyword matching.
-Each record’s covered_dimensions field (Section 3) provides the gold set, namely dimension IDs explicitly labeled by annotators during critique creation.
-All models generate English critiques; DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
-DCR is intended as a coarse diagnostic rather than a leaderboard metric.
-On a 50-sample validation set, DCR correlates with human-annotated dimension counts (Pearson $r=0.82$).
-Expert review further estimates 78% precision for keyword-detected hits.
-To test robustness to adversarial manipulation, we applied keyword stuffing; this increases surface-level hits but yields only 31% DCR, indicating limited susceptibility.
-For semantic validation, we report agreement rates between keyword detections and two independent checks at the sample level: (i) embedding-based similarity (cosine $\geq 0.7$) and (ii) NLI entailment, yielding 86% and 82% overall agreement, respectively.
-As expected, agreement is lower for L3–L5 (81% / 76%) due to partial mentions and context-dependent terminology.
-Full protocol details (thresholds, dictionaries, and length-controlled analyses) appear in Appendix D.
-6.2 Dimension Coverage Results
+Each record’s covered_dimensions field provides the gold set, namely dimension IDs explicitly labelled by annotators during critique elaboration.
+All our models generate critiques in English language, then DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
+A complete description of the DCR validation protocol can be found in Appendix D.
+5.2 DCR Results
 Model L1–L2 L3–L5 $\Delta_{L}$ DCR
@@
 GLM-4V-Flash 78.4 40.7 37.7 58.2
-Table 7: Layer-Gap Diagnostic results.
-$\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit.
-All models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5).
-These values should be interpreted as a reference rather than state-of-the-art results because DCR is a coarse proxy and the pilot uses a limited prompt configuration.
-N=336 balanced-pilot subset (7 cultures); bootstrap 95% CI half-width $\approx$ 4.8 pp.
-All models exhibit a 25–40 percentage-point layer gap ($\Delta_{\text{L}}=\text{DCR}_{\text{L1--L2}}-\text{DCR}_{\text{L3--L5}}$), confirming Vulca-Bench captures cultural understanding beyond visual perception.
-The 14-point DCR spread (72.4%–58.2%) demonstrates discriminative power across model capabilities.
-Critically, model orderings on the balanced-pilot subset (N=336, 7 cultures) correlate strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C).
+Table 6: DCR results. $\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit. Results on N=336 balanced-pilot subset (7 cultures).
+As observed in Table 6 (Figure 3, left), all models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5) i.e., $\Delta_{\text{L}}$, confirming that Vulca-Bench captures cultural understanding beyond just visual perception.
+The 14-point DCR spread (72.4%–58.2%) also highlights cultural-coverage differences across model capabilities.
+Interestingly, model performance on the balanced-pilot subset (N=336, 7 cultures) also correlates strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C).
 The balanced-pilot subset enables controlled, equal-weighted cross-cultural comparisons at low evaluation cost.
-The full corpus, by contrast, supports aggregate benchmarking with narrower uncertainty estimates due to the larger sample size.
+On the other hand, the full corpus supports aggregated benchmarking with narrower uncertainty estimates due to the larger sample size.
 5.3 Implications
-These results confirm our hypothesis that cultural understanding is fundamentally hierarchical.
-The consistent 25–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability tier from visual perception.
+These results confirm our initial hypothesis that cultural understanding is fundamentally hierarchical.
+The consistent 31–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability from visual perception.
 This finding implies that architectural improvements targeting cultural understanding should focus specifically on higher-layer reasoning rather than general visual capability.
-We release and validate DCR (78% precision; resistant to adversarial manipulation) as a reproducible, dataset-level diagnostic check signal for dimension coverage.
 5.4 Error Analysis: VLM Failures on L3–L5
 Beyond aggregate metrics, Vulca-Bench enables fine-grained error diagnosis. We present three representative failure patterns exposing how VLMs struggle with higher-order cultural reasoning.
-We identify three recurring error patterns (detailed examples in Appendix H): (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c.
+We identify three recurring error patterns: (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c.
 vanitas to 16th-c.
@@
 Mughal miniatures).
+Surface-level terminology.
+GPT-4o analysing a Song Dynasty landscape scroll generates the following: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations.
+Expert comment identifies three visual strategies instead: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs. dian), and strategic negative space (liubai).
+Historical anachronism.
+Claude-Sonnet-4.5 on a 16th-century Dutch market scene: “This vanitas painting conveys Protestant anxieties about mortality… characteristic of 17th-century Dutch Golden Age memento mori.” The VLM applies 17th-century conventions to a 1560s pronkstilleven emphasizing Renaissance prosperity rather than later Calvinist meditation.
+Cultural conflation.
+Gemini-2.5-Pro identifies a Safavid Persian miniature as: “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
 These errors demonstrate that L3–L5 layers require specialised cultural knowledge currently lacking in VLMs.
-Vulca-Bench’s 195 dimensions provide diagnostic precision to categorise such failures systematically.
+Vulca-Bench’s 225 dimensions provide diagnostic precision to categorise such failures systematically.
 5.5 Few-Shot Learning with Expert Critiques
 We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs).
-We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars (Table 8).
-Contrary to expectations, few-shot prompting yields decreased performance across all tested VLMs.
+We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars.
+Contrary to expectations, as observed in Table 7 and Figure 3 (right), few-shot prompting yields decreased performance across all tested VLMs.
 DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$).
@@
 Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
-We hypothesise three contributing factors: (1) attention dilution—longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting—exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity—exposure to expert critique structures may constrain flexibility.
-The stark model-dependent variation suggests more capable models (Claude) better resist these effects through superior context management, while others (DeepSeek-VL2) are more susceptible.
+We hypothesise three contributing factors: (1) attention dilution, longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting, exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity, exposure to expert critique structures may constrain flexibility.
+The noticeable model-dependent variation suggests more capable models (i.e., Claude) better resist these effects through superior context management, while others (e.g., DeepSeek-VL2) are more susceptible.
 […]
@@
 Significance: *** $p<0.001$.
-7 Release and Licensing
-The complete dataset is available under CC BY 4.0 in the supplementary materials.
-The release package includes: metadata JSON (7,410 pairs with bilingual critiques and dimension annotations), all images (optimised to $\leq$3.75MB), 225 dimension definitions, synonym dictionaries, evaluation scripts, and annotation guidelines.
-We release the full evaluation package and provide cached model outputs to support reproducibility.
-Using cached responses, the pilot experiments can be replicated at under USD 20 in API costs at the time of writing (see reproduce_pilot.py for the exact configuration).
-All artworks are from public museum collections (public domain or CC0); each record includes source_institution and accession_number for provenance (see Appendix I for culture distribution visualization).
 6 Conclusion
-We present Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, operationalised through a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions.
-The dataset addresses critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
-Vulca-Bench enables systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception.
-The complete dataset (images, metadata, critiques, dimension definitions, and evaluation/annotation tools) is available under CC BY 4.0 in the supplementary materials.
-We emphasise that DCR serves as a coarse-grained diagnostic rather than a leaderboard metric.
-Vulca-Bench enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset.
-The dataset complements, not replaces, human cultural expertise.
+We presented Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, annotated considering a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions.
+The dataset addresses three critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
+Vulca-Bench, therefore enables a systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception, focusing on critical interpretive aspects such as historical context or philosophical aesthetics.
+Vulca-Bench also enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset.
+We also report first-of-its-kind results on multi-layered cultural coverage, highlighting that while state-of-the-art VLMs tend to correctly assess visual and technical aspects, they struggle with deeper concepts such as cultural symbolism, historical context and philosophical aesthetics.
 Limitations
@@
 Second, L5 (philosophical aesthetics) inherently involves interpretive judgment, with higher reviewer correction rates at L5 (3.8%) compared to L1–L2 (0.5%) in our 100-sample audit, confirming higher subjectivity at philosophical levels.
-Third, the dataset reflects existing museum digitisation, which may underrepresent certain periods or genres within each tradition.
+Third, the dataset reflects existing museum digitisation, which may under-represent certain periods or genres within each tradition.
 Fourth, our Cultural Symmetry evidence (layer difficulty $r=0.96$, IAA parity) demonstrates structural comparability but not formal psychometric calibration; cross-cultural item-response validation remains future work.
@@
 We encourage community efforts to develop parallel corpora in additional languages.
-Finally, DCR serves as a coarse diagnostic rather than a precision instrument, designed for dataset utility validation rather than definitive model ranking.
+Finally, DCR serves as a coarse diagnostic rather than a thorough evaluation, designed for dataset utility validation rather than definitive model ranking.
 Ethical Considerations.
@@
 Appendix A Dimension Definitions
-Full dimension definitions for each culture are provided in the supplementary materials. Example dimensions:
+Full dimension definitions for each culture are available at https://github.com/yha9806/VULCA-Bench. Example dimensions:
 Chinese (30 dimensions).
@@
 Full 7,410 8 225 Complete corpus
-Eval 7,214 7 195 VLM evaluation (excl. Hermitage)
+Eval 7,214 7 195 VLM eval (excl. Hermitage)
 Gold 294 6 165 Expert references
 Human 450 6 Tier III calibration
-Balanced 384 8 225 Fairness analysis (48/culture)
-Balanced-Pilot 336 7 195 Pilot evaluation (excl. Hermitage)
-Table 9: Dataset variants.
-C=cultures, D=dimensions.
-8-culture: Full, Balanced (include Hermitage).
-7-culture: Eval, Balanced-Pilot (exclude Hermitage).
-6-culture: Gold, Human (exclude Mural, Hermitage due to insufficient gold annotations).
-Hash b8a34e5f* verifies single source.
+Balanced 384 8 225 Fairness (48/culture)
+Balanced-Pilot 336 7 195 Pilot (excl. Hermitage)
+Table 8: Dataset variants. C=cultures, D=dimensions. 8-culture: Full, Balanced (incl. Hermitage). 7-culture: Eval, Balanced-Pilot (excl. Hermitage). 6-culture: Gold, Human (excl. Mural, Hermitage).
 Appendix C Full-Corpus Ordering Consistency
@@
 Dimensions covered: JP_L1_D2, JP_L2_D1, JP_L3_D4, JP_L5_D2
-Appendix G Open-Source Embedding Replication
-To ensure reproducibility without proprietary API dependencies, we replicated the semantic alignment analysis (Section 6) using the open-source BGE-large-en-v1.5 model.
-[^2] Xiao et al., 2023.
-C-Pack: Packaged Resources To Advance General Chinese Embedding.
-https://arxiv.org/abs/2309.07597
-Layer OpenAI BGE $\Delta$
-L1–L2 94% 92% –2%
-L3–L5 81% 79% –2%
-Overall 86% 84% –2%
-Table 11: Semantic alignment comparison: OpenAI text-embedding-3-small vs BGE-large-en-v1.5 (open-source).
-Both models show the same trend: L1–L2 >> L3–L5, with 2% lower absolute agreement for BGE.
-The L1–L2 vs L3–L5 gap is preserved (13% for OpenAI, 13% for BGE), confirming that our DCR validation findings are not dependent on proprietary embeddings.
-Reproducibility: BGE-large-en-v1.5 is available on Hugging Face (BAAI/bge-large-en-v1.5) under MIT license. We release the replication script (eval/semantic_align_bge.py) with the dataset.
-Appendix H Error Pattern Examples
-Surface-level terminology.
-GPT-4o analysing a Song Dynasty landscape scroll produces: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations.
-Expert correction identifies three visual strategies: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs.
-dian), and strategic negative space (liubai).
-Historical anachronism.
-Claude-Sonnet-4.5 on a 16th-century Dutch market scene: “This vanitas painting conveys Protestant anxieties about mortality… characteristic of 17th-century Dutch Golden Age memento mori.” The VLM applies 17th-century conventions to a 1560s pronkstilleven emphasizing Renaissance prosperity rather than later Calvinist meditation.
-Cultural conflation.
-Gemini-2.5-Pro identifies a Safavid Persian miniature as “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
-Appendix I Culture Distribution Visualization
+Appendix G Culture Distribution Visualization
 Figure 4: Dataset Distribution.

### v3

@@
 Our pilot results indicate that higher-layer reasoning (L3–L5) is consistently more challenging than visual and technical analysis (L1–L2).
-The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 on GitHub [^1] https://github.com/yha9806/VULCA-Bench and Hugging Face [^2] https://huggingface.co/datasets/harryHURRY/vulca-bench.
+The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 at https://github.com/yha9806/VULCA-Bench.
 VULCA-Bench: A Multicultural Vision-Language Benchmark
 for Evaluating Cultural Understanding
-Haorui Yu1, Ramon Ruiz-Dolz2, and Diji Yang3, and Hang He4, and Fengrui Zhang5, and Qiufeng Yi6 1DJCAD, University of Dundee, United Kingdom 2Faculty of Science, Engineering and Business, University of Dundee, United Kingdom 3 University of California, Santa Cruz, USA 4East China Normal University, China 5Nanjing University, China 6School of Computer Science, University of Birmingham, United Kingdom
+Haorui Yu1, Diji Yang2,3, Hang He4, Fengrui Zhang5, and Qiufeng Yi6 1DJCAD, University of Dundee, United Kingdom 2University of California, Santa Cruz, USA 3Analogy AI 4East China Normal University, China 5Nanjing University, China 6Department of Mechanical Engineering, School of Engineering, University of Birmingham
 1 Introduction
 Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks.
-For instance, on object hallucination benchmarks like POPE Li et al.
-(2023) where leading models such as InstructBLIP reach 88.7% accuracy, and on visual question answering such as VQAv2 Goyal et al.
-(2017) where best models exceed 75% accuracy.
+On object hallucination benchmarks like POPE Li et al.
+(2023), leading models such as InstructBLIP reach 88.7% accuracy, and on visual question answering (VQAv2 Goyal et al.
+(2017)), frontier models exceed 75% accuracy.
 However, these benchmarks predominantly assess visual perception, which involves recognising objects, describing scenes, and answering factual questions.
@@
 (2025).
-We define cultural understanding by using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics).
-Higher layers require progressively deeper cultural knowledge (see Section 3 for complete definitions).
+We define cultural understanding using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics).
+Higher layers require progressively deeper cultural knowledge (Section 3 provides full definitions).
 Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3), the artist’s lineage within the “Four Gentlemen” (四君子) tradition (L4), or aesthetic principles such as 气韵生动 (qiyun shengdong, “spirit resonance”) and 意境 (yijing, “artistic conception”) Jing (2023) that define Chinese painting philosophy (L5).
 This example illustrates that cultural understanding is not a single capability but a progression across interpretive levels, a progression that existing benchmarks fail to measure.
-Recent cultural AI datasets (e.g., CulturalBench Chiu et al.
+Recent cultural AI datasets (CulturalBench Chiu et al.
 (2025), CultureAtlas Fung et al.
@@
 Initial work on Chinese painting critique Yu et al.
-(2025a) introduced a quantitative VLM evaluation method using 163 expert commentaries, revealing significant VLM-expert divergence.
+(2025a) introduced quantitative VLM evaluation using 163 expert commentaries, revealing significant VLM-expert divergence.
 Complementary probing Yu et al.
-(2025b) found VLMs rely on “symbolic shortcuts”, excelling on Western festivals but failing on under-represented traditions.
+(2025b) found VLMs rely on “symbolic shortcuts,” excelling on Western festivals but failing on underrepresented traditions.
 However, these studies focused on single cultures; systematic cross-cultural evaluation remains absent.
-Furthermore, existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
-This paper introduces Vulca-Bench (Vision-Understanding and Language-based Cultural Assessment), a comprehensive multicultural art critique benchmark addressing these gaps.
-Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and expert-validated factual accuracy, annotated using our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size).
+Existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
+We introduce Vulca-Bench (Vision-Understanding-Language-Culture Assessment), a comprehensive multicultural art critique benchmark addressing these gaps.
+Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and 98% cultural-fact accuracy, operationalised through our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size).
 All critiques meet expert-quality standards (Chinese $\geq$150 characters, English $\geq$100 words) with cultural specialist validation.
+The complete dataset is available in the supplementary materials.
 Vulca-Bench enables systematic evaluation of VLMs’ hierarchical cultural understanding, identification of L1–L5 performance gaps, equal-weighted cultural fairness probing on the balanced-pilot subset (N=336, 7 cultures), and development of culturally grounded VLM architectures.
 Figure 1 showcases representative artworks from each tradition.
-This paper therefore makes three contributions.
+This paper makes three contributions.
 (1) We present Vulca-Bench, a multicultural art-critique benchmark enabling evaluation of cultural interpretation beyond surface-level perception, comprising 7,410 image–critique pairs across eight traditions with 225 expert-defined dimensions.
@@
 (2024) covers 27 dimensions with 24K multiple-choice questions, and POPE Li et al.
-(2023) probes object hallucination through binary yes/no questions.
-However, none address cultural understanding in-depth.
+(2023) probes object hallucination through binary yes/no questions—yet none address cultural understanding.
 Specialised benchmarks such as MathVista Lu et al.
@@
 (2024) adopt QA formats rather than generative critique; GIMMICK Schneider et al.
-(2025) reveals Western biases across 144 countries employing recognition tasks; WuMKG Wan et al.
-(2024) provides a Chinese painting knowledge graph (104K entities) useful for information retrieval rather than generative evaluation; and CulTi Yuan et al.
+(2025) reveals Western biases across 144 countries but employs recognition tasks; WuMKG Wan et al.
+(2024) provides a Chinese painting knowledge graph (104K entities) for retrieval, not generative evaluation; and CulTi Yuan et al.
 (2026) offers image-text pairs for heritage retrieval without hierarchical cultural dimensions.
@@
 (2021) focuses on affective responses; and Zhang et al.
+Zhang et al.
 (2024) survey computational approaches for traditional Chinese painting based on Xie He’s Six Principles, focusing primarily on visual and technical analysis.
@@
 Vulca-Bench addresses all three gaps through cross-cultural extension to eight traditions.
-Unlike QA-based cultural benchmarks that test factual recall or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5).
-Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (e.g., layer-wise analysis or dimension coverage scoring) unavailable in existing datasets.
+Unlike QA-based cultural benchmarks that test factual recall, or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5).
+Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (layer-wise drop analysis, dimension coverage scoring) unavailable in existing datasets.
 3 Dataset Construction
@@
 This design reduces Western-centric methodological bias while acknowledging that museum availability and expert access naturally produce imbalanced sample sizes.
-Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage, whose European court art overlaps with Western).
-We recommend using balanced variants for cross-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
+Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage).
+We recommend using balanced variants for per-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
 We operationalise Cultural Symmetry through three pillars.
@@
 We therefore set this threshold empirically, based on the observation that lower coverage correlated strongly with missing philosophical or historical analysis.
-Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates distributed proportionally across cultures).
+Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates, distributed proportionally across cultures).
 Beyond structural parity, we validate measurement alignment using three criteria.
-First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures.
-All cultures show the same pattern: L1 easiest, L5 hardest ($r=0.96$ cross-culture correlation).
-Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.79$ for L1–L2 and $0.71$ for L3–L5, with cross-culture standard deviation below 0.04.
+First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures; the resulting five-point layer profiles are highly similar (mean pairwise Pearson correlation across cultures $r=0.96$).
+Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.80$ for L1–L2 and $0.72$ for L3–L5, with cross-culture standard deviation below 0.04.
 Third, each culture’s L5 captures its deepest interpretive challenge, supporting functional equivalence.
@@
 3.2 Cultural Understanding Framework
-We operationalise cultural understanding through our five-layer framework reflecting increasing interpretive complexity.
-While L1–L2 require visual observation, L3–L5 demand progressively deeper cultural and contextual knowledge.
+Our five-layer hierarchy is grounded in Panofsky’s (1939) iconological method, which distinguishes pre-iconographic description (recognising pure forms and objects), iconographic analysis (identifying conventional themes and symbols), and iconological interpretation (grasping intrinsic cultural and philosophical meaning).
+We operationalise this ascending hierarchy—from perception through convention to interpretation—as five levels that isolate empirically separable VLM competencies.
+L1 (visual perception) captures the sensory recognition at the core of Panofsky’s pre-iconographic stratum, while L2 (technical analysis) foregrounds medium and technique identification that practical observation presupposes yet that demands domain training.
+L3 (cultural symbolism) targets the conventional subject-matter decoding of iconographic analysis, while L4 (historical context) isolates art-historical attribution, which Panofsky distributes across his second and third strata.
+L5 (philosophical aesthetics) corresponds to iconological interpretation.
+L1–L2 require visual observation; L3–L5 demand progressively deeper cultural knowledge that VLMs typically lack.
 Layer Description
@@
 Each culture adapts L1–L5 to its indigenous aesthetic vocabulary, yielding 225 dimensions total.
-Dimension counts (25–30) reflect tradition complexity.
-For example, Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境).
-On the other hand, Western cultures cover 25 dimensions that emphasise more formal analysis (chiaroscuro, perspective).
-In Vulca-Bench, all critiques must cover at least 70% of their respective culture-specific dimensions, ensuring substantive L3–L5 engagement.
+The dimensions were developed through an iterative process: for each tradition, we surveyed established analytical frameworks in the art-historical literature (e.g., Panofsky’s iconological method for Western art, Xie He’s Six Principles for Chinese painting, rasa theory for Indian art), then performed pilot tagging on a sample of expert critiques to identify recurring analytical categories.
+The resulting codebook was refined through two revision rounds with domain-trained annotators.
+Dimension counts (25–30) reflect tradition complexity: Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境); Western 25 dimensions emphasise formal analysis (chiaroscuro, perspective).
+All critiques must cover $\geq$70% of culture-specific dimensions, ensuring substantive L3–L5 engagement.
 Culture Dims Pairs Key Concepts
@@
 Chinese artworks derive from the Palace Museum, Shanghai Museum, and Metropolitan Museum; Western works from the Metropolitan Museum, Louvre, and Hermitage Museum; Japanese from Tokyo and Kyoto National Museums; Korean from the National Museum of Korea; Islamic from the Museum of Islamic Art Doha and Metropolitan Museum; Indian from the National Museum Delhi and British Museum; and Hermitage works from the State Hermitage Museum collection (Russian and European court art, religious iconography).
-All images are released under CC BY 4.0 at https://github.com/yha9806/VULCA-Bench and HuggingFace: https://huggingface.co/datasets/harryHURRY/vulca-bench.
+All images are released in the supplementary materials.
 3.4 Bilingual Critique Annotation
@@
 Third, expert availability: our annotation team (Table 3) comprises bilingual specialists with deep cultural expertise, enabling controlled drafting, translation, and review within a single protocol.
-This design reflects practical constraints rather than linguistic preference, thus extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in the Limitations section.
+This design reflects practical constraints rather than linguistic preference; extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in Limitations.
 Culture Ann. Rev. Background
@@
 Indian 2 1 PhD South Asian art
-Table 3: Annotation team by culture. Ann.=annotators, Rev.=reviewers. All have 10+ years specialisation. Reviewers resolve disagreements and ensure cultural accuracy.
+Table 3: Annotation team by culture.
+Ann.=annotators, Rev.=reviewers.
+All have 10+ years specialisation.
+Reviewers resolve disagreements and ensure cultural accuracy.
+Mural critiques are produced by the Chinese team (Buddhist art specialists) and Hermitage critiques by the Western team (European court art specialists), given cultural overlap with their parent traditions.
 Expert annotators generate critiques following the L1–L5 framework.
@@
 A reviewer validates each critique on dimension coverage ($\geq$70%), cultural accuracy, and bilingual consistency.
-Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources.
-The complete scheme of our annotation guidelines can be found in Appendix E and representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
+Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources; unresolved cases (<<2%) escalate to senior expert panel.
+Full annotation guidelines appear in Appendix E; representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
 We measure inter-annotator agreement on a 100-sample subset (Table 4).
@@
 Korean follows a modified protocol due to the scarcity of qualified Korean art-history specialists with dual fluency.
-Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with a 100% expert review by a senior scholar.
-The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA.
+Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with 100% expert review by a senior scholar.
+The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA; we report Korean separately for methodological transparency.
 Culture N L1–L2 $\kappa$ L3–L5 $\kappa$ ICC
@@
 Korean uses a different protocol: 1 primary annotator (PhD Korean studies) + 100% expert review.
+The reported $\kappa$ is annotator–reviewer agreement, not independent dual-annotation IAA; we report it separately for quality transparency but do not include it in pooled statistics.
+3.5 Quality Assurance
+Multi-stage validation ensures data quality.
+Each record contains an explicit dimension checklist through its covered_dimensions field, listing dimension IDs (e.g., ["CN_L1_D1", "CN_L3_D2", ...]) explicitly labeled by annotators during critique creation rather than derived from text post-hoc, providing auditable gold-standard dimension sets.
+Automated checks enforce the $\geq$70% dimension coverage threshold (21/30 for CN, 18/25 for WE, etc.) based on covered_dimensions length.
+We removed all exact duplicates after deep audit.
+Image-critique verification cross-validates artist/title/period against image content, achieving >>99% metadata match.
+All pairs have both Chinese ($\geq$150 characters) and English critiques, ensuring bilingual completeness.
+Finally, all images are compressed to $\leq$3.75MB for VLM API compatibility.
 4 Dataset Statistics
 Vulca-Bench comprises 7,410 matched image-critique pairs drawn from 11,964 total images across 8 cultural traditions, annotated with 225 culture-specific dimensions.
-Table 5 summarises key metrics.
-The average Chinese critique length is 450 characters, while English critiques have an average of 280 words.
-Bilingual coverage was complete with all samples having EN-CN critiques, and all duplicates were excluded.
+Table 5 summarises key metrics: average Chinese critique length is 450 characters (English: 280 words), with 100% bilingual completeness, zero duplicates, >99% metadata verification, and 98% cultural-fact accuracy.
 All pairs meet the $\geq$70% dimension coverage threshold.
@@
 […]
-Figure 2 reveals layer-wise coverage patterns.
-While L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding.
+Figure 2 reveals layer-wise coverage patterns: L1–L2 (visual/technical) achieve near-complete coverage ($\geq$93%), while L3–L5 (cultural/philosophical) show progressive decline to 70–89%, reflecting the interpretive complexity of deeper cultural understanding.
 Chinese art shows highest L5 coverage (78%), consistent with its rich philosophical aesthetics tradition (qiyun, yijing).
-Importantly, each culture’s critiques are authored by annotators with specialized expertise in that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
-L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapaneseKoreanIslamicIndianMuralHermitage
-Figure 2: Layer-wise dimension coverage across all right cultures. L1–L2 achieve $\geq$93%; L3–L5 decline to 70–89%, reflecting interpretive complexity.
+Importantly, each culture’s critiques are authored by native expert annotators from that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
+L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapanese
+Figure 2: Layer-wise dimension coverage across all eight cultures. L1–L2 achieve $\geq$93%; L3–L5 decline to 70–89%, reflecting interpretive complexity.
+5 Data Quality Validation
+We validate Vulca-Bench through a three-phase audit protocol.
+Verification ensures metadata match (artist, title, period align with image content), visual fidelity (described elements match the actual image), and cultural fact check (art-historical claims verified against authoritative sources).
+Phase 1 (automated) performs MD5 deduplication, bilingual completeness, and image validation.
+Phase 2 (stratified) constitutes the 450 human-scored subset.
+KR: 107 (full set), JP/IS/IN: 50 each, CN/WE: 100 each.
+Phase 3 (expert) validates 100 samples stratified by culture and layer group (14–17 per culture, balanced L1–L2 vs L3–L5) against museum catalogues and art-historical databases (e.g., Grove Art Online, Oxford Art Online).
+This yields an estimated cultural-fact accuracy of 98% with a Wilson CI of [93.0%, 99.8%]; we emphasise this is a sample-based estimate rather than a census of all 7,410 records.
+Reproducibility: Stratification uses a fixed random seed (42) with proportional allocation by culture size.
+Per-sample audit outcomes (pass/fail/corrected) are logged in audit_log_p3.csv (supplementary), and full reviewer assignments are documented in Appendix E.
+Table 6 reports quality metrics with Wilson confidence intervals: 100% bilingual completeness, 0% duplicates, 100% image availability (100-sample audit), and 100% dimension coverage ($\geq$70% threshold).
+To address sample-size imbalance concerns, we provide a balanced subset (N=48 per culture $\times$ 7, total 336 pairs).
+Multi-seed stability tests (10 random seeds) confirm robustness: model orderings are stable across resamples (standard deviation of rank positions, $\sigma_{\text{rank}}\leq 1.5$).
+Inter-culture variation also remains low (standard deviation of per-culture mean DCR across seeds, $\sigma\leq 0.05$).
+The L1–L2 versus L3–L5 layer-gap pattern is preserved in every resample, indicating that the observed trends are not artefacts of a particular sample selection.
+Metric Val 95% CI
+Bilingual complete 100% [99.9%, 100%]
+Duplicate rate 0% [0%, 0.1%]
+Image available 100% [96.3%, 100%]
+Dim $\geq$70% 100% [99.9%, 100%]
+Table 6: Quality metrics with Wilson CI. 100-sample audit.
+Vulca-Bench supports multiple research applications.
+The 7,410 expert critiques with L1–L5 structure enable cultural VLM fine-tuning, while 225 dimensions provide probing targets for interpretability research.
+The balanced subset (N=336 for pilot, 48/culture $\times$ 7; N=384 available for full evaluation with 8 cultures) enables equal-weighted cultural fairness probing, and bilingual critiques support cultural knowledge retrieval and RAG augmentation.
+[^1] Hermitage excluded from pilot due to late inclusion; full 8-culture evaluation (N=384) available in released dataset.
 6 Pilot Evaluation
-Gemini-2.5-ProQwen3-VL-235BGPT-4oClaude-Sonnet-4.5GLM-4V-Flash02020404060608080100100Coverage (%)(a) Layer-wise CoverageL1–L2 (Visual/Technical)L3–L5 (Cultural/Philosophical)DeepSeek-VL2GPT-4oClaude-OpusClaude-Sonnet05050100100Coverage (%)(b) Few-shot Prompting Results0-shot1-shot3-shot
-Figure 3: VLM diagnosis reveals systematic cultural understanding gaps.
-(a) Layer-wise coverage: all models exhibit 31–40 percentage-point gaps between L1–L2 and L3–L5, confirming cultural/philosophical layers are harder than visual/technical analysis.
-(b) Few-shot prompting results: DeepSeek-VL2 and GPT-4o show significant degradation from 0-shot to 3-shot; Claude models remain robust.
 To demonstrate Vulca-Bench’s utility as a cultural VLM benchmark, we evaluate 5 representative VLMs on a stratified sample.
-This pilot evaluation is aimed at showcasing dataset utility and does not represent a definitive benchmark to compare against.
-For that purpose, we use Dimension Coverage Rate (DCR) to verify that VLM outputs engage with culture-specific concepts and to highlight the usefulness of Vulca-Bench, specifically focusing on cultural coverage rather than on a complete assessment of the VLM-generated critiques.
+This pilot evaluation establishes dataset utility and diagnostic capability only; it does not constitute a definitive evaluation framework.
+We use Dimension Coverage Rate (DCR) as a dataset-level diagnostic check, verifying that VLM outputs engage with culture-specific concepts, rather than as a leaderboard metric for model ranking.
 6.1 Setup
-We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash.
-Temperature was set to 0.7 and max_tokens to 2048.
-We then use DCR as a measure to test whether model-generated critiques engage with culture-specific concepts.
+We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash (temperature=0.7, max_tokens=2048; full configuration in supplementary).
+We use Dimension Coverage Rate (DCR) as a diagnostic to test whether model-generated critiques engage with culture-specific concepts.
 For a critique $c$ from culture $k$:
@@
 where $\mathcal{D}_{k}$ is the culture-specific dimension set (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$) and $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ denotes the subset of dimensions detected via keyword matching.
-Each record’s covered_dimensions field provides the gold set, namely dimension IDs explicitly labelled by annotators during critique elaboration.
-All our models generate critiques in English language, then DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
-A complete description of the DCR validation protocol can be found in Appendix D.
-5.2 DCR Results
+Each record’s covered_dimensions field (Section 3) provides the gold set, namely dimension IDs explicitly labeled by annotators during critique creation.
+All models generate English critiques; DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
+On a 50-sample validation set, DCR correlates with human-annotated dimension counts (Pearson $r=0.82$).
+Expert review further estimates 78% precision for keyword-detected hits.
+To test robustness to adversarial manipulation, we applied keyword stuffing; this increases surface-level hits but yields only 31% DCR, indicating limited susceptibility.
+For semantic validation, we report agreement rates between keyword detections and two independent checks at the sample level: (i) embedding-based similarity (cosine $\geq 0.7$) and (ii) NLI entailment, yielding 86% and 82% overall agreement, respectively.
+As expected, agreement is lower for L3–L5 (81% / 76%) due to partial mentions and context-dependent terminology.
+Full protocol details (thresholds, dictionaries, and length-controlled analyses) appear in Appendix D.
+6.2 Dimension Coverage Results
 Model L1–L2 L3–L5 $\Delta_{L}$ DCR
@@
 GLM-4V-Flash 78.4 40.7 37.7 58.2
-Table 6: DCR results. $\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit. Results on N=336 balanced-pilot subset (7 cultures).
-As observed in Table 6 (Figure 3, left), all models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5) i.e., $\Delta_{\text{L}}$, confirming that Vulca-Bench captures cultural understanding beyond just visual perception.
-The 14-point DCR spread (72.4%–58.2%) also highlights cultural-coverage differences across model capabilities.
-Interestingly, model performance on the balanced-pilot subset (N=336, 7 cultures) also correlates strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C).
+Table 7: Layer-Gap Diagnostic results.
+$\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit.
+All models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5).
+These values should be interpreted as a reference rather than state-of-the-art results due to a limited prompt configuration.
+N=336 balanced-pilot subset (7 cultures); bootstrap 95% CI half-width $\approx$ 4.8 pp.
+All models exhibit a 25–40 percentage-point layer gap ($\Delta_{\text{L}}=\text{DCR}_{\text{L1--L2}}-\text{DCR}_{\text{L3--L5}}$), confirming Vulca-Bench captures cultural understanding beyond visual perception.
+The 14-point DCR spread (72.4%–58.2%) demonstrates discriminative power across model capabilities.
+Critically, model orderings on the balanced-pilot subset (N=336, 7 cultures) correlate strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C).
 The balanced-pilot subset enables controlled, equal-weighted cross-cultural comparisons at low evaluation cost.
-On the other hand, the full corpus supports aggregated benchmarking with narrower uncertainty estimates due to the larger sample size.
+The full corpus, by contrast, supports aggregate benchmarking with narrower uncertainty estimates due to the larger sample size.
 6.3 Implications
-These results confirm our initial hypothesis that cultural understanding is fundamentally hierarchical.
-The consistent 31–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability from visual perception.
+These results confirm our hypothesis that cultural understanding is fundamentally hierarchical.
+The consistent 25–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability tier from visual perception.
 This finding implies that architectural improvements targeting cultural understanding should focus specifically on higher-layer reasoning rather than general visual capability.
+We release and validate DCR (78% precision; resistant to adversarial manipulation) as a reproducible, dataset-level diagnostic check signal for dimension coverage.
 6.4 Error Analysis: VLM Failures on L3–L5
 Beyond aggregate metrics, Vulca-Bench enables fine-grained error diagnosis. We present three representative failure patterns exposing how VLMs struggle with higher-order cultural reasoning.
-We identify three recurring error patterns: (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c.
+We identify three recurring error patterns (detailed examples in Appendix H): (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c.
 vanitas to 16th-c.
@@
 Mughal miniatures).
-Surface-level terminology.
-GPT-4o analysing a Song Dynasty landscape scroll generates the following: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations.
-Expert comment identifies three visual strategies instead: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs. dian), and strategic negative space (liubai).
-Historical anachronism.
-Claude-Sonnet-4.5 on a 16th-century Dutch market scene: “This vanitas painting conveys Protestant anxieties about mortality… characteristic of 17th-century Dutch Golden Age memento mori.” The VLM applies 17th-century conventions to a 1560s pronkstilleven emphasizing Renaissance prosperity rather than later Calvinist meditation.
-Cultural conflation.
-Gemini-2.5-Pro identifies a Safavid Persian miniature as: “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
 These errors demonstrate that L3–L5 layers require specialised cultural knowledge currently lacking in VLMs.
 Vulca-Bench’s 225 dimensions provide diagnostic precision to categorise such failures systematically.
-5.5 Few-Shot Learning with Expert Critiques
-We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs).
-We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars.
-Contrary to expectations, as observed in Table 7 and Figure 3 (right), few-shot prompting yields decreased performance across all tested VLMs.
-DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$).
-GPT-4o follows with $15.5\%$ decline ($p<0.001$).
-Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
-We hypothesise three contributing factors: (1) attention dilution, longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting, exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity, exposure to expert critique structures may constrain flexibility.
-The noticeable model-dependent variation suggests more capable models (i.e., Claude) better resist these effects through superior context management, while others (e.g., DeepSeek-VL2) are more susceptible.
-[…]
-Model N 0-shot 1-shot 3-shot $\Delta$ $p$
-DeepSeek-VL2 362 0.985 0.819 0.578 -41.3% <<0.001
-GPT-4o 338 0.998 0.967 0.843 -15.5% <<0.001
-Claude-Opus 183 1.000 0.998 0.986 -1.4% 0.074
-Claude-Sonnet 145 1.000 0.992 0.985 -1.5% 0.124
-Table 7: Few-shot prompting performance (DCR) across 4 VLMs on Chinese and Western art (NN=1,028).
-DeepSeek-VL2 and GPT-4o show significant degradation; Claude models remain stable.
-$\Delta$: relative change from 0-shot to 3-shot.
-Significance: *** $p<0.001$.
+We additionally evaluate few-shot prompting with culture-matched expert critiques across four VLMs (NN=1,028; Appendix J), finding that naive few-shot elicitation decreases performance for most models, with DeepSeek-VL2 showing $41.3\%$ degradation while Claude models remain stable ($2\%$).
+7 Release and Licensing
+The complete Vulca-Bench corpus—images, bilingual critiques, dimension definitions, and evaluation tools—is released under CC BY 4.0.222Code and evaluation tools: https://github.com/yha9806/VULCA-Bench.
+Dataset: https://huggingface.co/datasets/harryHURRY/vulca-bench.
+All artworks derive from public museum collections (public domain or CC0) with full provenance metadata.
+We provide cached model outputs and evaluation scripts to enable full replication of pilot experiments; technical details appear in the supplementary README.
 8 Conclusion
-We presented Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, annotated considering a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions.
-The dataset addresses three critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
-Vulca-Bench, therefore enables a systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception, focusing on critical interpretive aspects such as historical context or philosophical aesthetics.
-Vulca-Bench also enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset.
-We also report first-of-its-kind results on multi-layered cultural coverage, highlighting that while state-of-the-art VLMs tend to correctly assess visual and technical aspects, they struggle with deeper concepts such as cultural symbolism, historical context and philosophical aesthetics.
+We present Vulca-Bench, a multicultural art critique benchmark comprising 7,410 matched image–critique pairs across 8 traditions, operationalised through a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions.
+The dataset addresses critical gaps in existing VLM benchmarks: hierarchical cultural understanding beyond surface-level perception, non-Western aesthetic frameworks with equal methodological treatment, and expert-quality bilingual critique data.
+Pilot experiments on five VLMs confirm that cultural understanding is hierarchically structured: all models exhibit a 25–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5), with error analysis revealing surface-level terminology use, historical anachronism, and cultural conflation as recurring failure modes.
+These findings demonstrate that Vulca-Bench provides diagnostic precision that existing benchmarks lack.
+The dataset supports research into culturally aware AI systems, including cultural VLM fine-tuning, interpretability probing via the 225 dimensions, and cross-cultural fairness evaluation via the balanced subset.
+The dataset complements, not replaces, human cultural expertise.
+The complete corpus is released under CC BY 4.0 (§7).
 Limitations
@@
 Second, L5 (philosophical aesthetics) inherently involves interpretive judgment, with higher reviewer correction rates at L5 (3.8%) compared to L1–L2 (0.5%) in our 100-sample audit, confirming higher subjectivity at philosophical levels.
-Third, the dataset reflects existing museum digitisation, which may under-represent certain periods or genres within each tradition.
+Third, the dataset reflects existing museum digitisation, which may underrepresent certain periods or genres within each tradition.
 Fourth, our Cultural Symmetry evidence (layer difficulty $r=0.96$, IAA parity) demonstrates structural comparability but not formal psychometric calibration; cross-cultural item-response validation remains future work.
@@
 Finally, DCR serves as a coarse diagnostic rather than a thorough evaluation, designed for dataset utility validation rather than definitive model ranking.
+A comprehensive tri-tier evaluation framework with judge-based scoring and human calibration is developed in the companion study (Yu et al., 2026).
 Ethical Considerations.
@@
 We mitigate this through source attribution (all dataset JSON records include provenance metadata linking critiques to Vulca-Bench), detection research (we encourage using our bilingual critiques as positive examples for training AI-generated content detectors), and usage terms (CC BY 4.0 license requires attribution, creating audit trails for derivative works).
+Acknowledgments
+We thank Ramon Ruiz-Dolz for initial discussions on project framing, early-stage conceptual feedback, and manuscript polishing during the formative phase of this work.
 Appendix A Dimension Definitions
-Full dimension definitions for each culture are available at https://github.com/yha9806/VULCA-Bench. Example dimensions:
+Full dimension definitions for each culture are provided in the supplementary materials. Example dimensions:
 Chinese (30 dimensions).
@@
 Table 8 clarifies the relationship between dataset variants.
-Key boundary: Full corpus and Balanced include all 8 cultures (with Hermitage); Eval uses 7 cultures (Hermitage excluded); Gold/Human subsets use 6 cultures (Mural and Hermitage excluded due to insufficient gold annotations for reliable Tier III calibration).
+Key boundary: Full corpus and Balanced include all 8 cultures (with Hermitage); Eval uses 7 cultures (Hermitage excluded); Gold/Human subsets use 6 cultures (Mural and Hermitage excluded due to insufficient gold annotations for reliable Tier III calibration; see Yu et al., 2026).
 Variant Pairs C D Usage
@@
 Balanced-Pilot 336 7 195 Pilot (excl. Hermitage)
-Table 8: Dataset variants. C=cultures, D=dimensions. 8-culture: Full, Balanced (incl. Hermitage). 7-culture: Eval, Balanced-Pilot (excl. Hermitage). 6-culture: Gold, Human (excl. Mural, Hermitage).
+Table 8: Dataset variants.
+C=cultures, D=dimensions.
+8-culture: Full, Balanced (incl.
+Hermitage).
+7-culture: Eval, Balanced-Pilot (excl.
+Hermitage).
+6-culture: Gold, Human (excl.
+Mural, Hermitage).
+Hash b8a34e5f* verifies single source.
 Appendix C Full-Corpus Ordering Consistency
@@
 Dimensions covered: JP_L1_D2, JP_L2_D1, JP_L3_D4, JP_L5_D2
-Appendix G Culture Distribution Visualization
+Appendix G Open-Source Embedding Replication
+To ensure reproducibility without proprietary API dependencies, we replicated the semantic alignment analysis (Section 6) using the open-source BGE-large-en-v1.5 model.
+[^3] Xiao et al., 2023.
+C-Pack: Packaged Resources To Advance General Chinese Embedding.
+https://arxiv.org/abs/2309.07597
+Layer OpenAI BGE $\Delta$
+L1–L2 94% 92% –2%
+L3–L5 81% 79% –2%
+Overall 86% 84% –2%
+Table 10: Semantic alignment comparison: OpenAI text-embedding-3-small vs BGE-large-en-v1.5 (open-source).
+Both models show the same trend: L1–L2 >> L3–L5, with 2% lower absolute agreement for BGE.
+The L1–L2 vs L3–L5 gap is preserved (13% for OpenAI, 13% for BGE), confirming that our DCR validation findings are not dependent on proprietary embeddings.
+Reproducibility: BGE-large-en-v1.5 is available on Hugging Face (BAAI/bge-large-en-v1.5) under MIT license. We release the replication script (eval/semantic_align_bge.py) with the dataset.
+Appendix H Error Pattern Examples
+Surface-level terminology.
+GPT-4o analysing a Song Dynasty landscape scroll produces: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations.
+Expert correction identifies three visual strategies: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs.
+dian), and strategic negative space (liubai).
+Historical anachronism.
+Claude-Sonnet-4.5 on a 16th-century Dutch market scene: “This vanitas painting conveys Protestant anxieties about mortality… characteristic of 17th-century Dutch Golden Age memento mori.” The VLM applies 17th-century conventions to a 1560s pronkstilleven emphasizing Renaissance prosperity rather than later Calvinist meditation.
+Cultural conflation.
+Gemini-2.5-Pro identifies a Safavid Persian miniature as “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
+Appendix I Culture Distribution Visualization
 Figure 3: Dataset Distribution.
@@
 Western and Chinese dominate (82%); minority cultures maintain high quality through rigorous validation.
-Complete dataset is available under CC BY 4.0 in supplementary materials.
+Complete dataset is available in supplementary materials.
+Appendix J Few-Shot Learning Details
+We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs).
+We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus-4.5, and Claude-Sonnet-4.5, prepending $k$ culture-matched expert critiques as in-context exemplars (Table 11).
+Contrary to expectations, few-shot prompting yields decreased performance across all tested VLMs.
+DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$).
+GPT-4o follows with $15.5\%$ decline ($p<0.001$).
+Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
+[^4] Zero-shot DCR values in this experiment (0.985–1.000) are higher than the pilot DCR in Table 7 because the few-shot experiment uses only Chinese and Western art, whose well-established keyword dictionaries yield higher matching rates.
+The pilot includes all seven cultures, where minority traditions with sparser synonym coverage lower the aggregate.
+We hypothesise three contributing factors: (1) attention dilution—longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting—exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity—exposure to expert critique structures may constrain flexibility.
+The stark model-dependent variation suggests more capable models (Claude) better resist these effects through superior context management, while others (DeepSeek-VL2) are more susceptible.
+[…]
+Model N 0-shot 1-shot 3-shot $\Delta$ $p$
+DeepSeek-VL2 362 0.985 0.819 0.578 -41.3% <<0.001
+GPT-4o 338 0.998 0.967 0.843 -15.5% <<0.001
+Claude-Opus-4.5 183 1.000 0.998 0.986 -1.4% 0.074
+Claude-Sonnet-4.5 145 1.000 0.992 0.985 -1.5% 0.124
+Table 11: Few-shot prompting performance (DCR) across 4 VLMs on Chinese and Western art (NN=1,028).
+DeepSeek-VL2 and GPT-4o show significant degradation; Claude models remain stable.
+$\Delta$: relative change from 0-shot to 3-shot.
+Significance: *** $p<0.001$.